### v3

# It’s Not All Black and White: Degree of Truthfulness for Risk-Avoiding Agents

Abstract.
The classic notion of truthfulness requires that no agent has a profitable manipulation — an untruthful report that, for some combination of reports of the other agents, increases her utility. This strong notion implicitly assumes that the manipulating agent either knows what all other agents are going to report, or is willing to take the risk and act as-if she knows their reports.
Without knowledge of the others’ reports, most manipulations are risky – they might decrease the manipulator’s utility for some other combinations of reports by the other agents. Accordingly, a recent paper (Bu, Song and Tao, “On the existence of truthful fair cake cutting mechanisms”, Artificial Intelligence 319 (2023), 103904) suggests a relaxed notion, which we refer to as risk-avoiding truthfulness (RAT), which requires only that no agent can gain from a safe manipulation — one that is sometimes beneficial and never harmful.
Truthfulness and RAT are two extremes: the former considers manipulators with complete knowledge of others, whereas the latter considers manipulators with no knowledge at all. In reality, agents often know about some — but not all — of the other agents. This paper introduces the RAT-degree of a mechanism, defined as the smallest number of agents whose reports, if known, may allow another agent to safely manipulate, or $n$ if there is no such number. This notion interpolates between classic truthfulness (degree $n$) and RAT (degree at least $1$): a mechanism with a higher RAT-degree is harder to manipulate safely.
To illustrate the generality and applicability of this concept, we analyze the RAT-degree of prominent mechanisms across various social choice settings, including auctions, indivisible goods allocations, cake-cutting, voting, and two-sided matching.
conference: ; ;
1. Introduction
The Holy Grail of mechanism design is the truthful mechanism — a mechanism in which the (weakly) dominant strategy of each agent is truthfully reporting her type. But in most settings, there is provably no truthful mechanism that satisfies other desirable properties such as budget-balance, efficiency or fairness. Practical mechanisms are thus manipulable in the sense that some agent $a_{i}$ has a profitable manipulation – for some combination of reports by the other agents, agent $a_{i}$ can induce the mechanism to yield an outcome (strictly) better for her by reporting non-truthfully.
This notion of manipulability implicitly assumes that the manipulating agent either knows the reports made by all other agents, or is willing to take the risk and act as-if she knows their reports. Without knowledge of the others’ reports, most manipulations are risky – they might decrease the manipulator’s utility for some other combinations of reports by the other agents. In practice, many agents are risk-avoiding and will not manipulate in such cases. This highlights a gap between the standard definition and the nature of such agents.
To illustrate, consider a simple example in the context of voting. Under the Plurality rule, agents vote for their favorite candidate, and the candidate with the most votes wins. If an agent knows that her preferred candidate has no chance of winning, she may find it beneficial to vote for her second-choice candidate to prevent an even less preferred candidate from winning. However, if the agent lacks precise knowledge of the other votes and decides to vote for her second choice, it may backfire – she might inadvertently cause an outcome worse than if she had voted truthfully. For instance, if the other agents vote in a way that makes the agent the tie-breaker.
Indeed, various papers on cake-cutting (e.g. (brams2006better; BU2023Rat)), voting (e.g. (slinko2008nondictatorial; slinko2014ever; hazon2010complexity)) stable matching (e.g. regret2018Fernandez; chen2024regret) and coalition formation (waxman2021manipulation) studies truthfulness among such agents. In particular, BU2023Rat introduced a weaker notion of truthfulness, suitable for risk-avoiding agents, for cake-cutting. Their definition can be adapted to any problem as follows.
Let us first define a safe manipulation as a non-truthful report that may never harm the agent’s utility. Based on that, a mechanism is safely manipulable if some agent $a_{i}$ has a manipulation that is both profitable and safe; otherwise, the mechanism is Risk-Avoiding Truthful (RAT).
Standard truthfulness and RAT can be seen as two extremes with respect to safe-and-profitable manipulations: the former considers manipulators with complete knowledge of others, whereas the latter considers manipulators with no knowledge at all. In reality, agents often know about some — but not all — of the other agents.
This paper introduces the RAT-Degree — a new measurement that quantifies how robust a mechanism is to such safe-and-profitable manipulations. The RAT-Degree of a mechanism is an integer $d\in\{0,\ldots,n\}$, which represents — roughly — the smallest number of agents whose reports, if known, may allow another agent to safely manipulate; or $n$ if there is no such number. (See Section 3 for formal definition).
This measure allows us to position mechanisms along a spectrum. A higher degree implies that an agent has to work harder in order to collect the information required for a successful manipulation; therefore it is less likely that mechanism will be manipulated. On one end of the spectrum are truthful mechanisms – where no agent can safely manipulate even with complete knowledge of all the other agents. The RAT-degree of such mechanisms is $n$. While on the other end are mechanisms that are safely manipulable – no knowledge about other agents is required to safely manipulate. The RAT-degree of such mechanisms is 0.
Importantly, the RAT-degree is determined by the worst-case scenario for the mechanism designer, which corresponds to the best-case scenario for the manipulating agents. The way we measure the amount of knowledge is based on a general objective applicable to all social choice settings.
Contributions.
Our main contribution is the definition of the RAT-degree.
To illustrate the generality and usefulness of this concept, we selected several different social choice domains, and analyzed the RAT-degree of some prominent mechanisms in each domain. As our goal is mainly to illustrate the new concepts, we did not attempt to analyze all mechanisms and all special cases of each mechanism, but rather focused on some cases that allowed for a more simple analysis. To prove an upper bound on the RAT-degree, we need to show a profitable manipulation. However, in contrast to usual proofs of manipulability, we also have to analyze more carefully, how much knowledge on other agents is sufficient in order to guarantee the safety of the manipulation. This analysis gives more insight on the kind of manipulations possible in each mechanism, and on potential ways to avoid them. For clarity, we detail the results for each social choice setting in its corresponding section.
Organization.
Section 2 introduces the model and required definitions. Section 3 presents the definition of RAT-degree. Section 4 explores auctions for a single good. Section 5 examines indivisible goods allocations. Section 6 focuses on cake cutting. Section 7 addresses single-winner ranked voting. Section 8 considers two-sided matching. Section 9 concludes with some future work directions.
To enhance readability, we include only intuitive proof sketches in the main body of the paper. Complete and rigorous versions of all proofs are provided in the appendix.
1.1. Related Work
There is a vast body of work on truthfulness relaxations and alternative measurements of manipulability.
1.1.1. Truthfulness Relaxations
The large number of impossibility results have lead to extensive research into relaxations of truthfulness. A relaxed truthfulness notion usually focuses on a certain subset of all possible manipulations, which are considered more “likely”. It requires that none of the manipulations from this subset is profitable. Different relaxations consider different subsets of “likely” manipulations. Our approach can be combined with most of these ideas by adapting the degree concept to the relevant manipulation style.
RAT
Closest to our work is the recent paper by (BU2023Rat), which introduces the definition on which this paper is build upon: risk-avoiding truthfulness (RAT). The definition assumes that agents avoid risk - they will manipulate only when it is sometimes beneficial but never harmful. We first note that their original term for this concept is risk-averse truthfulness. However, since the definition assumes that agents completely avoid any element of risk, we adopt this new name, aiming to more accurately reflect this assumption.
We extend their work in two key directions. First, we generalize the definition from cake-cutting to any social choice problem. Second, we move beyond a binary classification of whether a mechanism is RAT, to a quantitative measure of its robustness to manipulation by such agents. Our new definition provides deeper insight of the mechanism’ robustness. In Section 6, we also analyze the mechanisms proposed in their paper alongside additional mechanisms for cake-cutting.
Their paper (in Section 5) provides an extensive comparison between RAT of related relaxation of truthfulness. For completeness, we include some of these comparisons here as well.
Dominating Manipulations with Partial Information.
conitzer2011dominating studied a closely related model in the context of voting, where the manipulator has only partial information about the preferences of the other (non-manipulating) voters and seeks a dominating manipulation, defined analogously to the profitable and safe manipulations we consider in this paper.
There are two main differences. First, their work focuses on the computational complexity of finding a dominating manipulation, while our work focuses on existence proofs. Second, their paper considers different information structures about the other agents — we assume that the manipulator knows exactly the preferences of some subset of voters, whereas they assume partial knowledge about all other voters. These two approaches can be combined: one could study the computational complexity under our information structure, or conversely, prove existence results under theirs.
As part of their analysis, they considered the special case in which the manipulator has no information about the other voters and showed that, for several voting rules, no dominating manipulation exists — in our terms, that the RAT-degree is positive. We extend some of these results in Section 7.
There has been a line of research following this paper — for example, Reijngoud2012 and endriss2016strategic. To the best of our knowledge, however, these studies focus specifically on voting settings, and the information structures they consider are tailored to that domain. In contrast, our measure of the number of known agents can be applied to any social choice problem, as illustrated later in the paper.
Veto Power.
Moulin1982vetoPower introduced the veto power for voting: a coalition of voters of size $t$ is said to have veto power $v(t)$ if it can prevent any set of $v(t)$ candidates from winning, regardless of the preferences of the remaining voters.
We conjectured that this notion could provide an upper bound on the RAT-degree: our hypothesis was that knowing that one of the candidate has no chance of winning is sufficient to enable a safe and profitable manipulation for a manipulator for some candidate. This indeed holds for Plurality and Veto rules (if the eliminated candidate is the manipulator’s favorite under Plurality, and the least favorite under Veto; see Section 7). Therefore, we expected that the minimal size of a coalition with positive veto power would serve as an upper bound on the RAT-degree. However, this claim appears to fail in general: eliminating a single candidate may not suffice for some manipulation to be both profitable and safe. A counterexample and several partial connections are presented in Section 7.3.
Maximin Strategy-Proofness
Another related relaxation of truthfulness, introduced by brams2006better, assumes the opposite extreme to the standard definition regarding the agents’ willingness to manipulate. In the standard definition, an agent is assumed to manipulate if for some combination of the other agents’ reports, it is beneficial. In contrast, this relaxation assumes that an agent will manipulate only if the manipulation is always beneficial — i.e., for any combination of the other agents’ reports.
This definition is not only weaker than the standard one but also weaker than RAT, as it assumes that agents manipulate in only a subset of the cases where RAT predicts manipulation. We believe that, in this sense, RAT provides a more realistic and balanced assumption. However, we note that a similar approach can be applied to this definition as well —- the degree to which a mechanism is robust to always-profitable manipulations (rather than to safe manipulations). The degree of Maximin Strategy-Proofness would probably be higher, as the manipulation is stronger.
NOM
troyan2020obvious introduce the notion of not-obvious manipulability (NOM), which focuses on obvious manipulation — informally, a manipulation that benefits the agent either in the worst case or in the best case. It presumes that agents are boundedly rational, and only consider extreme situations — best or worst cases. A mechanism is NOM if there are no obvious manipulations. This notion is a relaxation of truthfulness since the existence of an obvious manipulation imposes a stronger requirement than merely having a profitable one.
However, an obvious manipulation is not necessarily a safe manipulation, nor is a safe manipulation necessarily an obvious one (see Appendix D in (BU2023Rat) for more details). Therefore, NOM and RAT are independent notions.
Our approach can be applied to NOM as well.
Regret-Free Truth-telling (RFTT)
regret2018Fernandez proposes another relaxation of truthfulness, called regret-free truth-telling. This concept also considers the agent’s knowledge about other agents but does so in a different way. Specifically, a mechanism satisfies RFTT if no agent ever regrets telling the truth after observing the outcome – meaning that she could not have safely manipulated given that only the reports that are consistent with the observed outcome are possible. Note that the agent does not observe the actual reports of others but only the final outcome (the agents are assumed to know how the mechanism works).
An ex-post-profitable manipulation is always profitable, but the opposite is not true, as it is possible that the profiles in which a manipulation could be profitable are not consistent with the outcome. This means that RFTT does not imply RAT. A safe manipulation is always ex-post-safe, but the opposite is not true, as it is possible that the manipulation is harmful for some profiles that are inconsistent with the outcome. This means that RAT does not imply RFTT.
Different Types of Risk
slinko2008nondictatorial; slinko2014ever and hazon2010complexity study “safe manipulations” in voting. Their concept of safety is similar to ours, but allows a simultaneous manipulation by a coalition of voters. They focus on a different type of risk for the manipulators: the risk that too many or too few of them will perform the exact safe manipulation.
waxman2021manipulation study safe manipulations  [^1] They use “safe manipulation” for a manipulation that is both safe and profitable. on coalitional games in social networks. They examine the possible manipulations of agents with three different levels of knowledge on the social networks.
1.1.2. Alternative Measurements
Various measures can be used to quantify the manipulability of a mechanism. Below, we compare some existing approaches with our RAT-degree measure.
Computational Complexity
Even when an agent knows the reports of all other agents, it might be hard to compute a profitable manipulation. Mechanisms can be ranked according to the run-time complexity of this computation: mechanisms in which a profitable manipulation can be computed in polynomial time are arguably more manipulable than mechanisms in which this problem is NP-hard. This approach was pioneered by bartholdi1989computational; bartholdi1991single for voting, and applied in diverse social choice settings, e.g. coalitional games in social networks (waxman2021manipulation). See faliszewski2010ai; veselova2016computational for surveys.
Nowadays, with the advent of efficient SAT and CSP solvers, NP-hardness does not seem a very good defense against manipulation. Moreover, some empirical studies show that voting rules that are hard to manipulate in theory, may be easy to manipulate in practice (walsh2011hard). We claim that the main difficulty in manipulation is not the computational hardness, but rather the informational hardness — learning the other agents’ preferences.
Queries and Bits
Instead of counting the number of agents, we could count the number of bits that an agent needs to know in order to have a safe manipulation (this is similar in spirit to the concepts of communication complexity - e.g., (nisan2002communication; grigorieva2006communication; Communication2019Branzei; Babichenko2019communication) and compilation complexity - e.g., (chevaleyre2009compiling; xia2010compilation; karia2021compilation)). But this definition may be incomparable between different domains, as the bit length of the input is different. We could also measure the number of basic queries, rather than the number of agents. But then we have to define what “basic query” means, which may require a different definition for each setting. The number of agents is an objective measure that is relevant for all settings.
Probability of Having Profitable Manipulation.
Truthfulness requires that not even a single preference-profile allows a profitable manipulation. Mechanisms can be ranked according to the probability that such a profile exists – e.g., (barrot2017manipulation; lackner2018approval; lackner2023free). The probability of truthful mechanism is zero, and the lower the probability is, the more resistant the mechanism is to profitable manipulations.
A disadvantage of this approach is that it requires knowledge of the distribution over preference profiles; our RAT-degree does not require it. However, if we do have such information, it can be combined with our known-agents approach.
Incentive Ratio
Another common measure is the incentive ratio of a mechanism chen2011profitable, which describes the extent to which an agent can increase her utility by manipulating. For each agent, it considers the maximum ratio between the utility obtained by manipulating and the utility received by being truthful, given the same profile of the others. The incentive ratio is then defined as the maximum of these values across all agents.
This measure is widely used - e.g., (chen2022incentive; li2024bounding; cheng2022tight; cheng2019improved; bei2025incentive; tao2024fair). However, it is meaningful only when utility values have a clear numerical interpretation, such as in monetary settings. In contrast, our RAT-degree measure applies to any social choice setting, regardless of how utilities are represented.
Degree of Manipulability
aleskerov1999degree define four different indices of ”manipulability” of ranked voting rules, based on the fraction of profiles in which a profitable manipulation exists; the fraction of manipulations that are profitable; and the average and maximum gain per manipulation. As the indices are very hard to compute, they estimate them on 26 voting rules using computer experiments. These indices are useful under an assumption that all profiles are equally probable (“impartial culture”), or assuming a certain mapping between ranks of candidates and their values for the agent.
andersson2014budget; andersson2014least measure the degree of manipulability by counting the number of agents who have a profitable manipulation. They define a rule as minimally manipulable with respect to a set of acceptable mechanisms if for each preference profile, the number of agents with a profitable manipulation is minimal across all acceptable mechanisms.
2. Preliminaries
We consider a generic social choice setting, with a set of $n$ agents $N=\{a_{1},\ldots,a_{n}\}$, and a set of potential outcomes $X$. Each agent, $a_{i}\in N$, has preferences over the set of outcomes $X$, that can be described in one of two ways: (a) a linear ordering of the outcomes, or (b) a utility function from $X$ to $\mathbb{R}$. The set of all possible preferences for agent $a_{i}$ is denoted by $\mathcal{D}_{i}$, and is referred to as the agent’s domain. We denote the agent’s true preferences by $T_{i}\in\mathcal{D}_{i}$. Unless otherwise stated, when agent $a_{i}$ weakly prefers the outcome $x_{1}$ over $x_{2}$, it is denoted by $x_{1}\succeq_{i}x_{2}$; and when she strictly prefers $x_{1}$ over $x_{2}$, it is denoted by $x_{1}\succ_{i}x_{2}$.
A mechanism or rule is a function $f:\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{n}\to X$, which takes as input a list of reported preferences $P_{1},\ldots,P_{n}$ (which may differ from the true preferences), and returns the chosen outcome. In this paper, we focus on deterministic and single-valued mechanisms.
For any agent $a_{i}\in N$, we denote by $(P_{i},\mathbf{P}_{-i})$ the preference profile in which agent $a_{i}$ reports $P_{i}\in\mathcal{D}_{i}$, and the other agents report $\mathbf{P}_{-i}\in\mathcal{D}_{-i}$ (where $\mathcal{D}_{-i}:=\displaystyle\times_{j\in N\setminus\{i\}}\mathcal{D}_{j}$).
Truthfulness.
A manipulation for a mechanism $f$ and agent $a_{i}\in N$ is an untruthful report $P_{i}\in\mathcal{D}_{i}\setminus\{T_{i}\}$. A manipulation is profitable if there exists a set of preferences of the other agents for which it increases the manipulator’s utility:
(1) $\exists\mathbf{P}_{-i}\in\mathbf{\mathcal{D}}_{-i}:~f(P_{i},\mathbf{P}_{-i})\succ_{i}f(T_{i},\mathbf{P}_{-i})$
A mechanism $f$ is called manipulable if some agent $a_{i}$ has a profitable manipulation; otherwise $f$ is called truthful.
RAT
A manipulation is safe if it never harms the manipulator’s utility – it is weakly preferred over telling the truth for any possible preferences of the other agents:
(2) $\forall\mathbf{P}_{-i}\in\mathbf{\mathcal{D}}_{-i}:~f(P_{i},\mathbf{P}_{-i})\succeq_{i}f(T_{i},\mathbf{P}_{-i})$
A mechanism $f$ is called safely-manipulable if some agent $a_{i}$ has a manipulations that is profitable and safe; otherwise $f$ is called risk-avoiding truthful (RAT).
3. The RAT-Degree
[…]
Definition 3.1.
Given an agent $a_{i}$, a subset $K\subseteq N\setminus\{a_{i}\}$ and preferences for them $\mathbf{P}_{K}\in\mathcal{D}_{K}$:
A manipulation $P_{i}$ is called profitable for $a_{i}$ given $K$ and $\mathbf{P}_{K}$ if
[…]
A manipulation $P_{i}$ is called safe for $a_{i}$ given $K$ and $\mathbf{P}_{K}$ if
[…]
The previous two definitions are special cases of Definition 3.1: Equation 1 — which defines a profitable manipulation — is equivalent to $P_{i}$ being profitable given $\emptyset$; and Equation 2 — which defines a safe manipulation — is equivalent to $P_{i}$ being safe given $\emptyset$.
Definition 3.2.
A mechanism $f$ is called $k$-known-agents safely-manipulable if for some agent $a_{i}$, some subset $K\subseteq N\setminus\{a_{i}\}$ with $|K|=k$ and some preferences for them $\mathbf{P}_{K}$, there exists a manipulation $P_{i}$ that is both profitable and safe for $a_{i}$ given $K$ and $\mathbf{P}_{K}$.
Proposition 3.3.
prop:monotonicity Let $k\in\{0,\ldots,n-2\}$. If a mechanism is $k$-known-agents safely-manipulable, then it is also $(k+1)$-known-agents safely-manipulable.
LABEL:prop:monotonicity (proved in the appendix) justifies the following definition:
Definition 3.4.
The RAT-degree of a mechanism $f$ is the minimum $k$ for which the mechanism is $k$-known-agent safely-manipulable, or $n$ if there is no such $k$.
Intuitively, a mechanism with a higher RAT-degree is harder to manipulate, as a risk-avoiding agent would need to collect more information in order to find a safe manipulation.
Observation 3.5.
(a) A mechanism is truthful  if-and-only-if its RAT-degree is $n$.
(b) A mechanism is RAT if-and-only-if its RAT-degree is at least $1$.
Figure 1 illustrates the relation between classes of different RAT-degree.
Safely-Manipulable 11-KA Safely-Manipulable 22-KA Safely-Manipulable (n2)(n-2)-KA Safely-Manipulable (n1)(n-1)-KA Safely-Manipulable RAT-Degree: 0 11 22 $\cdots$ n1n-1 $n$ Truthful Risk-Avoiding Truthful (RAT)
Figure 1. Hierarchy of the Manipulability and Truthfulness Classes with respect to the RAT-Degree. The horizontal axis represents the RAT-Degree, from 0 (safely-manipulable) to $n$ (truthful). Labels above the axis correspond to Manipulability Classes, while labels below the axis correspond to Truthfulness Classes. KA stands for Known-Agents.
3.1. An Intuitive Point of View
Consider Table 1. We adopt the point of view of a particular agent $a_{i}$. The rows $T_{i},P^{1}_{i},\ldots,P^{3}_{i}$ correspond to the possible reports of agent $a_{i}$, where $T_{i}$ is the truthful report and the rest are potential manipulations. The columns $\mathbf{P}_{-i}^{1},\mathbf{P}_{-i}^{2},\ldots$ represent possible strategy profiles of the remaining n1n-1 agents. The values $x_{1},x_{2}\ldots$ indicate the utility of agent $a_{i}$ under each of these profiles when she reports truthfully.
When the risk-avoiding agent has no information (0-known-agents), a profitable-and-safe manipulation is a row in the table that represents an alternative report $P_{i}\neq T_{i}$, that (strictly) dominates $T_{i}$. That is, in each column, the outcome of the manipulation is at least as good as the truthful outcome, and in at least one column it is strictly better. In this example, $P^{1}_{i}$ satisfies this property.
When the risk-avoiding agent has more information ($k$-known-agents, when $k>0$), it is equivalent to considering a strict subset of the columns. For instance, suppose the agent can infer—based on the information she has over some $k$ other agents—that only the profiles $\mathbf{P}_{-i}^{3},\mathbf{P}_{-i}^{4},\mathbf{P}_{-i}^{5}$ are possible. Then $P^{2}_{i}$ is safe and profitable given this set. Notice that the utilities of agent $a_{i}$ in profiles not among $\mathbf{P}_{-i}^{3},\mathbf{P}_{-i}^{4},\mathbf{P}_{-i}^{5}$ are irrelevant, since $a_{i}$ considers them impossible.
Lastly, when the risk-avoiding agent has a full information ((n1)(n-1)-known-agents), she knows the exact strategy profile of the other agents, so it is equivalent to consider only one column in which the manipulation is profitable. $P^{3}_{i}$ in the table illustrates this type of manipulation.
$\mathbf{P}_{-i}^{1}$ $\mathbf{P}_{-i}^{2}$ $\mathbf{P}_{-i}^{3}$ $\mathbf{P}_{-i}^{4}$ $\mathbf{P}_{-i}^{5}$ $\mathbf{P}_{-i}^{6}$ $\mathbf{P}_{-i}^{7}$ $\ldots$
$T_{i}$ $x_{1}$ $x_{2}$ $x_{3}$ $x_{4}$ $x_{5}$ $x_{6}$ $x_{7}$ $\ldots$
[…]
Table 1. A Safe-And-Profitable Manipulation from an Agent Perspective. Dark-gray cells mean the value must be strictly higher, light-gray means it must be at least as high, and white means any value is allowed.
4. Auctions for a Single Good
We consider a seller owning a single good, and $n$ potential buyers (the agents). The true preferences $T_{i}$ of buyer $a_{i}$ are given by real values $v_{i}\geq 0$, representing her happiness from receiving the good. The reported preferences $P_{i}$ are the “bids” $b_{i}\geq 0$. A mechanism in this context has to determine the winner — the agent who will receive the good, and the price — how much the winner will pay. We assume that agents are quasi-linear – meaning their valuations can be interpreted in monetary units. Thus, the utility of the winning agent is the valuation minus the price; while the utility of the other agents is zero.
Results.
Table 2 provides a summary of our results. The two most well-known mechanisms in this context are first-price auction and second-price auctions. First-price auction maximizes the seller’s revenue when all buyers are truthful; but this assumption is, of course, unrealistic, as it is known to be manipulable. In fact, it is even safely-manipulable, so its RAT-degree is 0 (LABEL:thm:first-price-auction). We then prove that a first-price auction with a (fixed) positive discount has RAT-degree $1$.
Second-price auction is known to be truthful, so its RAT-degree is $n$. However, it has some important practical disadvantages (ausubel2006lovely). In particular, when buyers are risk-averse, the expected revenue of a second-price auction is lower than that of a first-price auction (nisan2007algorithmic); even when the buyers are risk-neutral, a risk-averse seller would prefer the revenue distribution of a first-price auction (krishna2009auction).
This raises the question of whether it is possible to combine the advantages of both auction types. Indeed, we prove that any auction that applies a weighted average between the first-price and the second-price [^2] The average price auction is mentioned as an exercise in (krishna2009auction) (Problem 2.4). achieves a RAT-degree of n1n-1, which is very close to being fully truthful (RAT-degree $n$). This implies that a manipulator agent would need to obtain information about all n1n-1 other agents to safely manipulate – which is a very challenging task. The seller’s revenue from such an auction is higher compared to the second-price auction, giving this mechanism a significant advantage in this context. This result opens the door to exploring new mechanisms that are not truthful but come very close it. Such mechanisms may enable desirable properties that are unattainable with truthful mechanisms.
Mechanism RAT-Degree
First-Price 0
First-Price With Positive Discount 11
Average Between First and Second Price n1n-1
Second-Price $n$
Table 2. Auctions for a Single Good: Summary of Results
4.1. First-Price Auction
In first-price auction, the agent who bids the highest price wins the good and pays her bid; the other agents get and pay nothing.
It in well-known that the first-price auction is not truthful. We start by proving that the situation is even worse: first-price auction is safely-manipulable, meaning that its RAT-degree is 0.
Theorem 4.1.
thm:first-price-auction First-price auction is safely manipulable (RAT-degree = 0).
Proof sketch.
A truthful agent always gets a utility of 0, whether he wins or loses. On the other hand, an agent who manipulates by bidding slightly less than $v_{i}$ gets a utility of at least 0 when he loses and a strictly positive utility when he wins. Thus, this manipulation is safe and profitable. ∎
4.2. First-Price With Positive Discount Auction
In first-price auction with discount, the agent $i$ with the highest bid $b_{i}$ wins the item and pays $(1-t)b_{i}$, where $t\in(0,1)$ is a constant. As before, the other agents get and pay nothing.
We prove that, although this minor change does not make the mechanism truthful, it increases the degree of trustfulness for risk-avoiding agents. However, it is still quite vulnerable to manipulation as knowing the strategy of one other agent might be sufficient to safely-manipulate it.
Theorem 4.2.
auction-with-discount The RAT-degree of the First-Price Auction with Discount is $1$.
To prove that the RAT-degree is $1$, we need to show that the mechanism is (a) not safely-manipulable, and (b) $1$-known-agents safely-manipulable. We prove each of these in a lemma.
Lemma 4.3.
First-Price Auction with Discount is not safely-manipulable.
Proof sketch.
Here, unlike in the first-price auction, whenever the agent wins the good, she gains positive utility. This implies that no manipulation can be safe. If she under-bids her value, she risks losing the item, which strictly decreases her utility. If she over-bids, she might end up paying more than necessary, reducing her utility without increasing her chances of winning. ∎
Lemma 4.4.
First-Price Auction with Discount is $1$-known-agents safely-manipulable.
Proof sketch.
Consider the case where the manipulator $a_{i}$ knows there is another agent $a_{j}$ bidding $b_{j}\in(v_{i},v_{i}/(1-t))$. When biding truthfully, she loses and gets zero utility. However, by bidding any value $b_{i}\in(b_{j},v_{i}/(1-t))$, she either wins and gains positive utility or loses and remains with zero utility. Thus it is safe and profitable. ∎
4.3. Average Between First and Second Price Auction
In the Average-First-Second-Price (AFSP) Auction, the agent $i$ with the highest bid $b_{i}$ wins the item and pays $wb_{i}+(1-w)b^{\max}_{-i}$, where $b^{\max}_{-i}:=\max_{j\neq i}b_{j}$ is the second-highest bid, and $w\in(0,1)$ is a fixed constant. That is, the price is a weighted average between the first price and the second price.
We show that this simple change makes a significantly difference – the RAT-degree increases to n1n-1. This means that a manipulator agent would need to obtain information about all other agents to safely manipulate the mechanism — a very challenging task in practice.
Theorem 4.5.
auction-average-price The RAT-degree of the Average-Price Auction is (n1)(n-1).
The theorem is proved using the following two lemmas.
Lemma 4.6.
The Average-Price Auction is not (n2)(n-2)-known-agents safely-manipulable.
Proof sketch.
Let $a_{i}$ be a potential manipulator, $b_{i}$ a potential manipulation, $K$ be a set of known agents, with |K|=n2|K|=n-2, and $\mathbf{b}_{K}$ be a vector that represents their bids. We prove that the only unknown-agent $a_{j}$ can make any manipulation either not profitable or not safe. We denote by $b^{\max}_{K}$ the maximum bid among the agents in $K$ and consider each of the six possible orderings of $v_{i}$, $b_{i}$ and $b^{\max}_{K}$. In two cases ($v_{i}<b_{i}<b^{\max}_{K}$ or $b_{i}<v_{i}<b^{\max}_{K}$) the manipulation is not profitable; in the other four cases, the manipulation is not safe. ∎
Lemma 4.7.
The Average-Price Auction is (n1)(n-1)-known-agents safely-manipulable.
Proof sketch.
Consider any combination of bids of the other agents in which all the bids are strictly smaller than $v_{i}$. Let $b^{\max}_{-i}$ be the highest bid among the other agents. Then any alternative bid $b_{i}\in(b^{\max}_{-i},v_{i})$ is a safe manipulation. ∎
Conclusion.
By choosing a high value for the parameter $w$, the Average-Price Auction becomes similar to the first-price auction, and therefore may attain a similar revenue in practice, but with better strategic properties. The average-price auction is sufficiently simple to test in practice; we find it very interesting to check how it fares in comparison to the more standard auction types.
5. Indivisible Goods Allocations
In this section, we consider several mechanisms to allocate $m$ indivisible goods $G=\{g_{1},\ldots,g_{m}\}$ among the $n$ agents. Here, the true preferences $T_{i}$ of agent $a_{i}$ are given by $m$ real values: $v_{i,\ell}\geq 0$ for any $g_{\ell}\in G$, representing her happiness from receiving the good $g_{\ell}$. The reported preferences $P_{i}$ are real values $r_{i,\ell}\geq 0$. We assume the agents have additive valuations over the goods. Given a bundle $S\subseteq G$, let $v_{i}(S)=\sum_{g_{\ell}\in S}v_{i,\ell}$ be agent $a_{i}$’s utility upon receiving the bundle $S$. A mechanism in this context gets $n$ (potentially untruthful) reports from all the agents and determines the allocation – a partition $(A_{1},\ldots,A_{n})$ of $G$, where $A_{i}$ is the bundle received by agent $a_{i}$.
Results.
Table 3 summarizes our results. We start by considering a simple mechanism, the utilitarian goods allocation – which assigns each good to the agent who reports the highest value for it. This mechanism always returns an allocation that maximizes the social welfare (i.e., sum of utilities) and is thus Pareto Optimal — there does not exist another allocation that all agents weakly prefer and at-least one strictly prefers. This mechanism is safely manipulable (RAT-degree = 0). We then show that the RAT-degree can be increased to $1$ by requiring normalization — the values reported by each agent are scaled such that the set of all items has the same value for all agents.
We then consider the famous round-robin mechanism. It is easy to compute (specifically, polynomial) and guarantees envy-freeness up to one good (EF1) — each agent weakly prefers their own bundle over each other agent’s bundle by the removal of at most one of the goods. amanatidis2017truthful prove that EF1 is incompatible with truthfulness for $m\geq 5$, which implies that the degree is at most n1n-1. We prove that the situation is much worse, the RAT-degree of round-robin is at most $1$. The proof relies on weak preferences (i.e., preferences that allow ties). In sharp contrast, we prove that when all agents have strict preferences, the RAT-degree is the best possible n1n-1.
Next, we turn to the Max-Nash-Welfare (MNW) rule, which selects an allocation that maximizes the product of agents’ valuations. This rule always outputs an allocation that is EF1 and Pareto optimal, and it also provides a good approximation to the maximin share (MMS) guarantee (see (caragiannis2019unreasonable) for more details). We examine three variants of the MNW mechanism, which differ in how they handle tie-breaking when multiple allocations achieve the same Nash welfare. We show that the RAT-degree of all these variants is at most $1$.
Lastly, we design a new mechanism that we call volatile priority. It satisfies EF1 and attains the best possible RAT-degree of n1n-1 for all types of preferences; but does not run in polynomial time. The main observation behind our approach is that, in many cases, there are multiple solutions (here, allocations) that satisfy the required fairness properties (here, EF1). We use a priority order over the agents, determined by their reported valuations, to select among these solutions. As a result, a manipulating agent who lacks information about all other agents’ valuations may unintentionally lower her priority by misreporting, and thus end up worse off. This makes manipulations risky.
This raises the following open question regarding the computational complexity:
Open Question 5.1.
Is there a polynomial time EF1 goods allocation rule with RAT-degree n1n-1?
A natural candidate is The Envy-Cycle Elimination Mechanism by lipton2004approximately.
Open Question 5.2.
What is the RAT-degree of The Envy-Cycle Elimination Mechanism by lipton2004approximately?
Mechanism RAT-Degree Properties
Utilitarian 0 Pareto Efficient
Normalized Utilitarian 11 Pareto Efficient
(w.r.t. the normalized values)
Max-Nash-Welfare $\leq 1$ $\leq$ EF1, Pareto Efficient
Round-Robin (for $n>m$) Any preferences 11 EF1, Polynomial-time
Strict preferences n1n-1
Volatile Priority* n1n-1 EF1
Table 3. Indivisible Goods Allocations: Summary of Results. (*) marks mechanisms introduced in this paper.
5.1. Utilitarian Goods Allocation
The utilitarian rule aims to maximize the sum of agents’ utilities. When agents have additive utilities, this goal can be achieved by assigning each good to an agent who reports the highest value for it. This section analyzes this mechanism.
We make the practical assumption that agents’ reports are bounded from above by some maximum possible value $r_{max}>0$.
We assume that in cases where multiple agents report the same highest value for some good, the mechanism employs some tie-breaking rule to allocate the good to one of them. However, the tie-breaking rule must operate independently for each good, meaning that the allocation for one good cannot depend on the tie-breaking outcomes of other goods.
Theorem 5.3.
prop:auction-knownagents The Utilitarian allocation rule is safely manipulable (RAT-degree = 0).
Proof sketch.
Manipulating by reporting the highest possible value, $r_{max}$, for all goods is both profitable and safe. It is profitable because if the maximum report among the other agents for a given good lies between the manipulator’s true value and their alternative bid, $r_{max}$, the manipulator’s utility strictly increases. It is safe because, in all other cases, the utility remains at least as high. ∎
5.2. Normalized Utilitarian Goods Allocation
In the normalized utilitarian allocation rule, the agents’ reports are first normalized such that each agent’s values sum to a given constant $V>0$. Then, each good is given to the agent with the highest normalized value. We focus on the case of at least three goods.
Theorem 5.4.
thm:normalized-utilitarian-goods For $m\geq 3$ goods, the RAT-degree of the Normalized Utilitarian allocation rule is $1$.
We prove LABEL:thm:normalized-utilitarian-goods using several lemmas that analyze different cases.
The first two lemmas prove that the rule is not safely-manipulable, so its RAT-degree is at least $1$. LABEL:claim:normalized-agent-who-likes-single-good addresses agents who value only one good positively, while LABEL:claim:normalized-agent-who-likes-at-least-two covers agents who value at least two goods positively.
Lemma 5.5.
claim:normalized-agent-who-likes-single-good An agent who values only one good positively cannot safely manipulate the Normalized Utilitarian allocation rule.
Proof sketch.
Due to the normalization requirement, any manipulation by such an agent involves reporting a lower value for the only good she values positively, while reporting a higher values for some goods she values at 0. This reduces her chances of winning her desired good and raises her chances of winning goods she values at zero, ultimately decreasing her utility. Thus, the manipulation is neither profitable nor safe. ∎
Lemma 5.6.
claim:normalized-agent-who-likes-at-least-two An agent who values positively at least two goods cannot safely manipulate the Normalized Utilitarian allocation rule.
Proof sketch.
Since values are normalized, any manipulation by such an agent must involve increasing the reported value of at least one good $g_{\mathrm{inc}}$ while decreasing the value of at least one other good $g_{\mathrm{dec}}$. We show that such a manipulation is not safe, by considering the case where all other agents report as follows: they assign a value of 0 to $g_{\mathrm{inc}}$, a value between the manipulator’s true and reported value for $g_{\mathrm{dec}}$, and a value slightly higher than the manipulator’s report for all other goods (it is possible to construct such reports that are non-negative and normalized).
With these reports, the manipulation causes the manipulator to lose the good $g_{\mathrm{dec}}$ which has a positive value for her, whereas she wins the good $g_{\mathrm{inc}}$ with or without the manipulation, and does not win any other good. Hence, the manipulation strictly decreases her total value. ∎
The last lemma shows that the RAT-degree is at most $1$, thus completing the proof of the theorem.
Lemma 5.7.
normalized-1-known With $m\geq 3$ goods, Normalized Utilitarian is $1$-known-agent safely-manipulable.
Proof sketch.
Consider a scenario where there is a known agent who reports 0 for some good $g$ that the manipulator wants and slightly more than the manipulator’s true values for all other goods. In this case, by telling the truth, the manipulator has no chance to win any good except $g$. Therefore, reporting a value of $V$ for $g$ and a value of 0 for all other goods is a safe manipulation.
The same manipulation is also profitable, since it is possible that the reports of all n2n-2 unknown agents for $g$ are larger than the manipulator’s true value and smaller than $V$. In such a case, the manipulation causes the manipulator to win $g$, which strictly increases her utility. ∎
5.3. Round-Robin Goods Allocation
In round-robin goods allocation, the agents are arranged according to some predetermined order $\pi$. There are $m$ rounds, corresponding to the number of goods. In each round, the agent whose turn it is (based on $\pi$) takes their most preferred item from the set of items that remain unallocated at that point. When there are multiple items that are equally most preferred, the agent breaks the tie according to a fixed item-priority ordering. Note that normalization is not important for Round-Robin, as it is not affected by scaling the valuations.
If there are at most $n$ items, then the rule is clearly truthful. Thus, we assume that there are $m\geq n+1$ items. Even the slight increment to $m=n+1$ makes a significant difference:
Theorem 5.8.
The RAT-degree of Round-Robin is:  (a) $n$ when $m\leq n$,  (b) $1$ when $m>n$;  and (c) n1n-1 when $m>n$ and all agents have strict preferences.
As (a) is trivial, we start by proving (b) using LABEL:RR-weak-prefs. The proof of the upper bound uses weak preferences; in contract, LABEL:thm:RR-strict-degree-n-1 proves that when all agents have strict preferences, the RAT-degree becomes n1n-1, yielding part (c).
Lemma 5.9.
RR-weak-prefs With $m\geq n+1$ items, the RAT-degree of Round-Robin is $1$.
Proof sketch.
We start by showing that Round-Robin is $1$-known-agent safely-manipulable. Consider the first turn of the agent who picks first, and the case where her two most preferred items are of equal value. If she knows that favorite item of the agent who picks right after is the item she planned to take first, she can safely manipulate by taking the other one instead. In some cases, this change allows her to obtain both items (since the second agent will then take a different one), and in all other cases she still ends up with the same utility as before. Hence, the manipulation is both profitable and safe.
Next, we prove that Round-Robin is not safely-manipulable. We first show that it is never safe to manipulate in a way that results in picking an item that is strictly less preferred than the best available one as it might reduce the manipulator’s utility. Thus, the only potentially profitable manipulations are those that swap between equally valued items. However, we show that if such a manipulation is profitable for some profile of the others’ preferences, we can construct a symmetric profile where the same manipulation is harmful. ∎
Lemma 5.10.
thm:RR-strict-degree-n-1 When all agents have strict preferences, the RAT-degree of Round-Robin is (n1)(n-1).
Proof sketch.
The challenging part is to show that it is not (n2)(n-2)-known-agents safely-manipulable. The idea is to track the set of available items after the manipulator’s turn — once under truthful reporting and once under misreporting — up to the turn of the unknown agent. We then show that there always exists a valuation for the unknown agent that causes the two bundles to become equal by the time the manipulator plays again. Since preferences are strict, the manipulator strictly loses in her first turn, and from that point on faces identical sets of available items. By an inductive argument, she can only lose from this point on. ∎
5.4. Max-Nash-Welfare item allocation
The Max-Nash-Welfare rule is defined by finding an allocation $(A_{1},\ldots,A_{n})$ that maximizes the Nash welfare $\prod_{i=1}^{n}v_{i}(A_{i})$. caragiannis2019unreasonable do not specify how to break ties when there is more than one allocation with the same Nash welfare. Here, we assume that ties are broken by a pre-specified agent priority. Specifically, when the optimal Nash welfare is positive and there are multiple allocations that maximize the Nash welfare, the mechanism outputs the allocation in the following lexicographical way: among all these allocations, consider those that maximize agent $a_{1}$’s valuation, and subject to this, maximize agent $a_{2}$’s valuation, and so on. When the optimal Nash welfare is 0 (i.e., at least one agent’s valuation is 0 for any allocation), there are many natural variants. Before describing them, let
[…]
and
$\mathcal{S}^{\max}=\{S\in\mathcal{S}\mid|S|\geq|S^{\prime}|\mbox{ for any }S^{\prime}\in\mathcal{S}\}.$
  • MNW1:
    Choose $S\in\mathcal{S}^{\max}$ with lexicographically optimal indices of agents (i.e., include $a_{1}\in S$ if there exists $S\in\mathcal{S}^{\max}$ that contains $a_{1}$, and, subject to this, include $a_{2}\in S$ if possible, and so on). Then find an allocation that maximizes $\prod_{a_{i}\in S}v_{i}(A_{i})$.
  • MNW2:
    Find the value of $\max_{S\in\mathcal{S}^{\max}}\max_{(A_{1},\ldots,A_{n})}\prod_{a_{i}\in S}v_{i}(A_{i})$. Find a lexicographically optimal $S$ where there exists an allocation such that $\prod_{a_{i}\in S}v_{i}(A_{i})$ equals this value. Output the corresponding allocation.
  • MNW3:
    Normalize the valuations of all agents such that $\sum_{j=1}^{m}v_{i,j}=1$ for each agent $a_{i}$. Then perform MNW2.
In words, all versions output an allocation that maximizes the number of agents with positive valuations. MNW1 performs the tie-breaking for the agent set first, and then proceeds to maximize the product (similarly to the definition in caragiannis2019unreasonable). MNW2 maximizes the product first, and then performs the tie-breaking among those agent sets that maximize the product. MNW3 is the variant of MNW2 by first normalizing the valuations (note that normalization has no effect on MNW1 because the selection of $S$ does not depend on the agents’ values).
Theorem 5.11.
For any $m\geq 2$ and any $n$, the RAT-degree of MNW2 is 0 regardless of tie-breaking.
The RAT-degrees of MNW1 and MNW3 with lexicographic agent-priority tie-breaking is at most $1$.
Given the superior fairness properties of the MNW rule, the following question is interesting.
Open Question 5.12.
Are there any variants of MNW with a RAT degree larger than 1?
5.5. Volatile Priority Goods Allocation: An EF1 Mechanism with RAT-degree n1n-1
In this section, we propose a new EF1 mechanism that has RAT-degree n1n-1. The mechanism starts by applying a given priority function $\Gamma$ (to be defined shortly), which takes as input a valuation profile $(v_{1},\ldots,v_{n})$ and outputs two agent indices, $i^{+}$ and $i^{-}$. Agent $a_{i^{+}}$ is referred to as the mechanism-favored agent, and agent $a_{i^{-}}$ as the mechanism-unfavored agent (the reason of both names will be clear soon). Let $\mathcal{X}^{\mathrm{EF1}}_{i^{-}}$ denote the set of all EF1 allocations $(A_{1},\ldots,A_{n})$ in which agent $a_{i^{-}}$ is not envied by any other agent; that is, for every agent $a_{i}$, we have $v_{i}(A_{i})\geq v_{i}(A_{i^{-}})$. Note that $\mathcal{X}^{\mathrm{EF1}}_{i^{-}}$ is non-empty: the allocation returned by the round-robin mechanism when $a_{i^{-}}$ is placed last in the picking order $\pi$ belongs to this set. The mechanism outputs the allocation in $\mathcal{X}^{\mathrm{EF1}}_{i^{-}}$ that maximizes the valuation of the mechanism-favored agent, i.e., it selects an allocation that maximizes $v_{i^{+}}(A_{i^{+}})$. When there are multiple maximizers, the rule breaks the tie in an arbitrary consistent way.
We show that the priority function $\Gamma$ can be constructed in a way that we refer to as volatile: lacking knowledge of even a single agent’s valuation makes manipulations risky. That is, for any agent $a_{i}$, any manipulation for $a_{i}$, any subset of (n2)(n-2)-known agents, and any valuations for them; there exists a valuation for the only remaining known-agent agent $a_{j}$ such that: (1) when $a_{i}$ reports truthfully, she is selected as the mechanism-favored agent; (2) when $a_{i}$ misreports, she becomes the mechanism-unfavored agent; and (3) this change in priority results in a strictly lower utility for $a_{i}$. A formal definition is provided in the appendix.
Any volatile priority rule $\Gamma$ can be used by our mechanism; to conclude its description we show that such a rule exists.
Proposition 5.13.
There exists a volatile priority rule $\Gamma$ for goods allocation.
Proof.
The rule $\Gamma$ first finds the maximum value among all agents and all goods: $v^{\ast}:=\max_{i\in[n],\ell\in[m]}v_{i,\ell}$. It then views the value $v^{\ast}$ as a binary string that encodes the following information:
  • the index $i$ of an agent $a_{i}$;
  • and the index $\ell$ of a good $g_{\ell}$
  • a non-negative integer $t$,
  • two non-negative integers $a,b$, between 0 and $n\choose 2$}.
We append 0’s as most significant bits to $v^{\ast}$ if the length of the binary string is not long enough to support the format of the encoding. If the encoding of $v^{\ast}$ is longer than the length enough for encoding the above-mentioned information, we take only the least significant bits in the amount required for the encoding.
The mechanism-favored agent $a_{i^{+}}$ and the mechanism-unfavored agent $a_{i^{-}}$ are then decided in the following way. Let $s\in\{0,1\}$ be the bit at the $t$-th position of the binary encoding of the value $v_{i}(g_{\ell})$.
Let $p:=(a\cdot s+b)\bmod{n\choose 2}$. Each value of $p$ corresponds to a pair of different agents $(i^{+},i^{-})$.
To see that $\Gamma$ is volatile, suppose $v_{i}$ and $v_{i}^{\prime}$ are different in the $t$-th bits of their binary encoding. We construct a value $v^{*}$ that encodes the integers $i,t,a,b$ where
  1. (1)
    the $t$-th bit of $v_{i}$ is $s$ and the $t$-th bit of $v_{i}^{\prime}$ is $s^{\prime}$ for $s\neq s^{\prime}$;
  2. (2)
    The pair $(i^{+},i^{-})$ corresponds to the integer $(a\cdot s+b)\bmod{n\choose 2}$.
  3. (3)
    The pair $(i^{+^{\prime}},i^{-^{\prime}})$ corresponds to the integer $(a\cdot s^{\prime}+b)\bmod{n\choose 2}$.
(1) can always be achieved by some encoding rule. To see (2) and (3) can always be achieved, assume $s=1$ and $s^{\prime}=0$ without loss of generality. We can then take $b:=$ the integer corresponding to the pair $(i^{+^{\prime}},i^{-^{\prime}})$, and a:=b+a:=-b+ the integer corresponding to the pair $(i^{+},i^{-})$, modulo $n\choose 2$}.
We then construct a valuation $v_{j}$ such that $v_{j,\ell^{\ast}}$ is the largest and is equal to $v^{*}$. In case $v^{*}$ is not large enough, we increase it as needed by adding most significant digits. ∎
Remark 5.14.
The proof of Proposition 5.13 requires that the mechanism does not normalize the valuations, nor place any upper bound on the reported values. Suppose there were an upper bound $V$ on the value. $V$ encodes some agent $i$, bit number $t$, and integers $a,b$. It is possible that these numbers give the highest priority to agent $i$. In that case, agent $i$ could manipulate by reporting the value $V$.
It is easy to see that the mechanism always outputs an EF1 allocation; we further prove that:
Theorem 5.15.
thm:volatile-good-alloc-degree-n-1 The RAT-degree of the Volatile Priority Goods Allocation is n1n-1.
Proof sketch.
For every profitable manipulation by $a_{i}$, and for every unknown agent $a_{j}$, the volatility of $\Gamma$ implies that, for some possible valuation $v_{j}$, a truthful report by $a_{i}$ leads to $a_{i}$ being the favored agent and $a_{j}$ being the unfavored agent, whereas the manipulation leads to $a_{i}$ being the unfavored agent and $a_{j}$ being the favored agent. We use this fact to prove that the manipulation may be harmful for $a_{i}$. ∎
Remark 5.16.
Our Volatile Priority algorithm does not run in polynomial time, as it requires to maximize the utility of a certain agent subject to EF1. which is an NP-hard problem (e.g., the proof in Appendix A.2 of barman2019fair can easily imply this). We do not know if a polynomial-time EF1 algorithm with a high RAT-degree exists.
In particular, an interesting direction is to to consider the round-robin algorithm but with a volatile priority order—that is, where the picking order depends on the reported valuations.
Open Question 5.17.
What is the RAT-degree of Volatile Priority Round Robin?
6. Cake Cutting
In this section, we study the cake cutting problem: the allocation of a divisible heterogeneous resource to $n$ agents. The cake cutting problem was proposed by Steinhaus48; Steinhaus49, and it is a widely studied subject in mathematics, computer science, economics, and political science.
In the cake cutting problem, the resource/cake is modeled as an interval $[0,1]$, and it is to be allocated among a set of $n$ agents $N=\{a_{1},\ldots,a_{n}\}$. An allocation is denoted by $(A_{1},\ldots,A_{n})$ where $A_{i}\subseteq[0,1]$ is the share allocated to agent $a_{i}$. We require that each $A_{i}$ is a union of finitely many closed non-intersecting intervals, and, for each pair of $i,j\in[n]$, $A_{i}$ and $A_{j}$ can only intersect at interval endpoints, i.e., the measure of $A_{i}\cap A_{j}$ is 0.
The true preferences $T_{i}$ of agent $a_{i}$ are given by a value density function $v_{i}:[0,1]\to\mathbb{R}_{\geq 0}$ that describes agent $a_{i}$’s preference over the cake. To enable succinct encoding of the value density function, we adopt the widely considered assumption that each $v_{i}$ is piecewise constant: there exist finitely many points $x_{i0},x_{i1},x_{i2},\ldots,x_{ik_{i}}$ with $0=x_{i0}<x_{i1}<x_{i2}<\cdots<x_{ik_{i}}=1$ such that $v_{i}$ is a constant on every interval $(x_{i\ell},x_{i(\ell+1)})$, $\ell=0,1,\ldots,k_{i}-1$. Given a subset $S\subseteq[0,1]$ that is a union of finitely many closed non-intersecting intervals, agent $a_{i}$’s value for receiving $S$ is then given by $V_{i}(S)=\int_{S}v_{i}(x)dx.$
Lastly, we define an additional notion, uniform segment, which will be used throughout this section. Given $n$ value density functions $v_{1},\ldots,v_{n}$ (that are piecewise constant by our assumptions), we identify the set of points of discontinuity for each $v_{i}$ and take the union of the $n$ sets. Sorting these points by ascending order, we let $x_{1},\ldots,x_{m-1}$ be all points of discontinuity for all the $n$ value density functions. Let $x_{0}=0$ and $x_{m}=1$. These points define $k$ intervals, $(x_{0},x_{1}),(x_{1},x_{2}),\ldots,(x_{m-1},x_{m})$, such that each $v_{i}$ is a constant on each of these intervals. We will call each of these intervals a uniform segment, and we will denote $X_{t}=(x_{t-1},x_{t})$ for each $t=1,\ldots,m$.
Results.
Our results are summarized in Table 4. We start by considering the utilitarian mechanism that outputs an allocation with the maximum social welfare. We show that the RAT-degree of this mechanism is 0. Similar as it is in the case of indivisible goods, we also consider the normalized variant of this mechanism, and we show that the RAT-degree is $1$. Both mechanisms are Pareto Efficient, they always returns an allocation that is Pareto Optimal — there is no other allocation that all agents weakly prefer, and at least one agent strictly prefers.
We then consider several fair mechanisms that have been considered by BU2023Rat. Their paper studies whether or not these mechanisms are risk-avoiding truthful (in our language, whether the RAT-degree is positive). Here we provide a more fine-grained view.
Each mechanism is guaranteed to return an allocation satisfying at least one of the following properties: Envy free (EF) — each agent weakly prefers their own bundle to the bundle of any other agent, Proportional — each agent receives at least $1/n$ of the total value of the cake, according to their own valuation; and Connected pieces — each agent receives a single connected piece.
We start by considering equal division mechanisms — where we evenly allocate each uniform segment $X_{t}$ to all agents. The output allocation is always envy-free and proportional. However, we show that the RAT-degree relies heavily on the order in which we allocate the pieces of each segment. Specifically, with fixed order (e.g., let agent $a_{1}$ get the left-most interval and agent $a_{n}$ get the right-most interval), BU2023Rat already proved that it is not even RAT, which means that its RAT-degree is 0. To avoid this type of manipulation, a different tie-breaking rule was considered by BU2023Rat (See Mechanism 3 in their paper). We prove that its RAT-degree is n1n-1.  tao2022existence shows that no EF mechanism can also be truthful, so n1n-1 is the best possible RAT-degree. However, this mechanism requires quite many cuts on the cake, and has a very poor performance in terms of pareto efficiency — it gives each agent a value of exactly $1/n$ and nothing more, which intuitively is the most inefficient allocation among the proportional ones.
We then focus on proportional mechanisms with connected pieces. BU2023Rat prove that the Dubins-Spanier’s Moving-Knife (dubins1961cut) is not RAT, which in our terms means that its RAT-degree is 0. There was hope for those proven by BU2023Rat to have a positive RAT-degree: Bu, Song, and Tao’s moving knife (BU2023Rat) (two variants - Mechanisms 4 and 5 in their paper); and Ortega and Segal-Halevi’s moving knife (ortega2022obvious). We show that they all have a very low RAT-degree of $1$. These results invoke the following question:
Open Question 6.1.
Is there a proportional connected cake-cutting rule with RAT-degree at least $2$?
A natural candidate to consider is the classic Even–Paz algorithm (even1984note); we conjecture that its RAT-degree is $\lceil n/2\rceil$, but the question remains open.
Open Question 6.2.
What is the RAT-degree of the Even-Paz algorithm?
Lastly, we adapt our volatile priority approach from Section 5.5, and propose a new mechanism that always returns a proportional and Pareto-efficient allocation, with the best possible RAT-degree of n1n-1. Unlike its counterpart, this mechanism runs in polynomial time.
Given this result, it is natural to ask if the fairness guarantee can be strengthened to envy-freeness. A compelling candidate is the mechanism that always outputs allocations with maximum Nash welfare — the product of agents utilities. It is well-known that such an allocation is EF and Pareto-efficient. However, computing its RAT-degree turns out to be very challenging for us. We conjecture the answer is n1n-1.
Open Question 6.3.
What is the RAT-degree of the maximum Nash welfare mechanism?
Mechanism RAT-Degree Properties
Utilitarian 0 Pareto Efficient
[…]
Normalized Utilitarian 11 Pareto Efficient
(w.r.t. the normalized values)
Bu, Song, and Tao’s Moving Knife 11 Proportional, Connected
(BU2023Rat) (Mechanism 4 and 5)
Ortega and Segal-Halevi’s Moving Knife 11 Proportional, Connected
Bu, Song, and Tao’s Equal Division n1n-1 Proportional, EF
(BU2023Rat) (Mechanism 3)
Volatile Priority* n1n-1 Proportional, Pareto Efficient
Table 4. Cake Cutting: Summary of Results. (*) indicates that the mechanism was proposed in this paper. Gray cells indicate indicate results proven by BU2023Rat.
6.1. Utilitarian Cake Cutting
It is easy to find an allocation that maximizes the social welfare: for each uniform segment $X_{t}$, allocate it to an agent $a_{i}$ with the maximum $v_{i}(X_{t})$. When multiple agents have equally largest value of $v_{i}(X_{t})$ on the segment $X_{t}$, we need to specify a tie-breaking rule. However, as we will see later, the choice of the tie-breaking rule does not affect the RAT-degree of the mechanism.
It is easy to see that, whatever the tie-breaking rule is, the maximum social welfare mechanism is safely manipulable. It is safe for an agent to report higher values on every uniform segment. For example, doubling the values on all uniform segments is clearly a safe manipulation.
Observation 6.4.
Utilitarian Cake-Cutting with any tie-breaking rule has RAT-degree 0.
We next consider the following variant of the maximum social welfare mechanism: first rescale each $v_{i}$ such that $V_{i}([0,1])=\int_{0}^{1}v_{i}(x)dx=1$, and then output the allocation with the maximum social welfare. We will show that the RAT-degree is $1$. The proof is similar to the one for indivisible items (LABEL:thm:normalized-utilitarian-goods) and is given in the appendix.
Theorem 6.5.
When there are at least three agents, Normalized Utilitarian Cake-Cutting with any tie-breaking rule has RAT-degree $1$.
6.2. Equal Division Cake Cutting
One natural envy-free mechanism is to evenly allocate each uniform segment $X_{t}$ to all agents. Specifically, each $X_{t}$ is partitioned into $n$ intervals of equal length, and each agent receives exactly one of them. It is easy to see that $V_{i}(A_{j})=\frac{1}{n}V_{i}([0,1])$ for any $i,j\in[n]$ under this allocation, so the allocation is envy-free and proportional.
To completely define the mechanism, we need to specify the order of evenly allocating each $X_{t}=(x_{t-1},x_{t})$ to the $n$ agents. A natural tie-breaking rule is to let agent $a_{1}$ get the left-most interval and agent $a_{n}$ get the right-most interval. Specifically, agent $a_{i}$ receives the $i$-th interval of $X_{t}$, which is $x_{t-1}+\frac{i-1}{n}(x_{t}-x_{t-1}),x_{t-1}+\frac{i}{n}(x_{t}-x_{t-1})]$. However, it was proved in BU2023Rat that the equal division mechanism under this ordering rule is safely-manipulable, i.e., its RAT-degree is 0. In particular, agent $a_{1}$, knowing that she will always receive the left-most interval in each $X_{t}$, can safely manipulate by deleting a point of discontinuity in her value density function if her value on the left-hand side of this point is higher.
To avoid this type of manipulation, a different ordering rule was considered by BU2023Rat (See Mechanism 3 in their paper): at the $t$-th segment, the $n$ equal-length subintervals of $X_{t}$ are allocated to the $n$ agents with the left-to-right order $a_{t},a_{t+1},\ldots,a_{n},a_{1},a_{2},\ldots,a_{t-1}$. By using this ordering rule, an agent does not know her position in the left-to-right order of $X_{t}$ without knowing others’ value density functions. Indeed, even if only one agent’s value density function is unknown, an agent cannot know the index $t$ of any segment $X_{t}$. This suggests that the mechanism has a RAT-degree of n1n-1.
Theorem 6.6.
Consider the mechanism that evenly partitions each uniform segment $X_{t}$ into $n$ equal-length subintervals and allocates these $n$ subintervals to the $n$ agents with the left-to-right order $a_{t},a_{t+1},\ldots,a_{n},a_{1},a_{2},\ldots,a_{t-1}$. It has RAT-degree n1n-1 and always outputs envy-free allocations.
Proof sketch.
Envy-freeness is trivial: for any $i,j\in[n]$, we have $V_{i}(A_{j})=\frac{1}{n}V_{i}([0,1])$. The general impossibility result in tao2022existence shows that no mechanism with the envy-freeness guarantee can be truthful, so the RAT-degree is at most n1n-1.
To show that the RAT-degree is exactly n1n-1, we show that, if even a single agent is not known to the manipulator, it is possible that this agent’s valuation adds discontinuity points in a way that the ordering in each uniform segment is unfavorable for the manipulator. ∎
Although the equal division mechanism with the above-mentioned carefully designed ordering rule is envy-free and has a high RAT-degree of n1n-1, it is undesirable in at least two aspects:
  1. (1)
    it requires quite many cuts on the cake by making n1n-1 cuts on each uniform segment; this is particularly undesirable if piecewise constant functions are used to approximate more general value density functions.
  2. (2)
    it is highly inefficient: each agent $a_{i}$’s utility is never more than her minimum proportionality requirement $\frac{1}{n}V_{i}([0,1])$;
We handle point (1) in Section 6.3 and point (2) in Section 6.4.
6.3. Connected Cake-Cutting
Researchers have been looking at allocations with connected pieces, i.e., allocations with only n1n-1 cuts on the cake. A well-known mechanism in this category is the moving-knife procedure, which always outputs proportional allocations. This mechanism was first proposed by dubins1961cut. It always returns a proportional connected allocation. Unfortunately, it was shown by BU2023Rat that Dubins and Spanier’s moving-knife procedure is safely-manipulable for some very subtle reasons. For clarity, we provide its description here.
Dubins and Spanier’s moving-knife procedure
Let $u_{i}=\frac{1}{n}V_{i}([0,1])$ be the value of agent $a_{i}$’s proportional share. In the first iteration, each agent $a_{i}$ marks a point $x_{i}^{(1)}$ on $[0,1]$ such that the interval $0,x_{i}^{(1)}]$ has value exactly $u_{i}$ to agent $a_{i}$. Take $x^{(1)}=\min_{i\in[n]}x_{i}^{(1)}$, and the agent $a_{i_{1}}$ with $x_{i_{1}}^{(1)}=x^{(1)}$ takes the piece $0,x^{(1)}]$ and leaves the game. In the second iteration, let each of the remaining n1n-1 agents $a_{i}$ marks a point $x_{i}^{(2)}$ on the cake such that $x^{(1)},x_{i}^{(2)}]$ has value exactly $u_{i}$. Take $x^{(2)}=\min_{i\in[n]\setminus\{i_{1}\}}x_{i}^{(2)}$, and the agent $a_{i_{2}}$ with $x_{i_{2}}^{(2)}=x^{(2)}$ takes the piece $x^{(1)},x^{(2)}]$ and leave the game. This is done iteratively until n1n-1 agents have left the game with their allocated pieces. Finally, the only remaining agent takes the remaining part of the cake. It is easy to see that each of the first n1n-1 agents receives exactly her proportional share, while the last agent receives weakly more than her proportional share; hence the procedure always returns a proportional allocation.
Remark 6.7.
Although such mechanism are usually described in an iterative interactive way that resembles an extensive-form game, we will consider the direct-revelation mechanisms in this paper, where the $n$ value density functions are reported to the mechanism at the beginning. In the above description of Dubins and Spanier’s moving-knife procedure, as well as its two variants mentioned later, by saying “asking an agent to mark a point”, we refer to that the mechanism computes such a point based on the reported value density function. In particular, we do not consider the scenario where agents can adaptively choose the next marks based on the allocations in the previous iterations.
BU2023Rat proposed a variant of the moving-knife procedure that is RAT. In addition, they showed that another variant of moving-knife procedure proposed by ortega2022obvious is also RAT. [^3] It should be noticed that, when $v_{i}$ is allowed to take 0 value, tie-breaking needs to be handled very properly to ensure RAT. See BU2023Rat for more details. Here, for simplicity, we assume $v_{i}(x)>0$ for each $i\in[n]$ and $x\in[0,1]$. Below, we will first describe both mechanisms and then show that both of them have RAT-degree $1$.
Ortega and Segal-Halevi’s moving knife procedure
The first iteration of Ortega and Segal-Halevi’s moving knife procedure is the same as it is in Dubins and Spanier’s. After that, the interval $x^{(1)},1]$ is then allocated recursively among the n1n-1 agents $n]\setminus\{a_{i_{1}}\}$. That is, in the second iteration, each agent $a_{i}$ marks a point $x_{i}^{(2)}$ such that the interval $x^{(1)},x_{i}^{(2)}]$ has value exactly $\frac{1}{n-1}V_{i}([x^{(1)},1])$ (instead of $\frac{1}{n}V_{1}([0,1])$ as it is in Dubins and Spanier’s moving-knife procedure). The remaining part is the same: the agent with the left-most mark takes the corresponding piece and leaves the game. After the second iteration, the remaining part of the cake is again recursively allocated to the remaining n2n-2 agents. This is continued until the entire cake is allocated.
Bu, Song, and Tao’s moving knife procedure
Each agent $a_{i}$ is asked to mark all the n1n-1 “equal-division-points” $x_{i}^{(1)},\ldots,x_{i}^{(n-1)}$ at the beginning such that $V_{i}([x_{i}^{(t-1)},x_{i}^{(t)}])=\frac{1}{n}V_{i}([0,1])$ for each $t=1,\ldots,n$, where we set $x_{i}^{(0)}=0$ and $x_{i}^{(n)}=1$. The remaining part is similar to Dubins and Spanier’s moving-knife procedure: in the first iteration, agent $i_{1}$ with the minimum $x_{i_{1}}^{(1)}$ takes $0,x_{i_{1}}^{(1)}]$ and leave the game; in the second iteration, agent $i_{2}$ with the minimum $x_{i_{2}}^{(2)}$ among the remaining n1n-1 agents takes $x_{i_{1}}^{(1)},x_{i_{2}}^{(2)}]$ and leave the game; and so on. The difference to Dubins and Spanier’s moving-knife procedure is that each $x_{i}^{(t)}$ is computed at the beginning, instead of depending on the position of the previous cut.
Theorem 6.8.
The RAT-degree of Ortega and Segal-Halevi’s moving knife procedure is $1$.
Proof.
It was proved in BU2023Rat that the mechanism is not 0-known-agent safely-manipulable. It remains to show that it is $1$-known-agents safely-manipulable. Suppose agent $a_{1}$’s value density function is uniform, $v_{1}(x)=1$ for $x\in[0,1]$, and agent $a_{1}$ knows that agent $a_{2}$ will report $v_{2}$ such that $v_{2}(x)=1$ for $x\in[1-\varepsilon,1]$ and $v_{2}(x)=0$ for $x\in[0,1-\varepsilon)$ for some very small $\varepsilon>0$ with $\varepsilon\ll\frac{1}{n}$. We show that the following $v_{1}^{\prime}$ is a safe manipulation.
$v_{1}^{\prime}(x)=\left\{\begin{array}[]{ll}1&x\in[0,\frac{n-2}{n}]\\ \frac{2}{n\varepsilon}&x\in[1-2\varepsilon,1-\varepsilon]\\ 0&\mbox{otherwise}\end{array}\right$.
Before we move on, note an important property of $v_{1}^{\prime}$: for any $t\leq\frac{n-2}{n}$, we have $V_{1}([t,1])=V_{1}^{\prime}([t,1])$.
Let $[a,b]$ be the piece received by agent $a_{1}$ when she reports $v_{1}$ truthfully. If $b\leq\frac{n-2}{n}$, the above-mentioned property implies that she will also receive exactly $[a,b]$ for reporting $v_{1}^{\prime}$. If $b>\frac{n-2}{n}$, then we know that agent $a_{1}$ is the (n1)(n-1)-th agent in the procedure. To see this, we have $V_{1}([a,b])\geq\frac{1}{n}([0,1])=\frac{1}{n}$ by the property of Ortega and Segal-Halevi’s moving knife procedure, and we also have $V_{1}([b,1])=1-b<\frac{2}{n}$. This implies $b$ cannot be the $1/k$ cut point of $[a,1]$ for $k\geq 3$. On the other hand, it is obvious that agent $a_{2}$ takes a piece after agent $a_{1}$. Thus, by the time agent $a_{1}$ takes $[a,b]$, the only remaining agent is agent $a_{2}$.
Since there are exactly two remaining agents in the game before agent $a_{1}$ takes $[a,b]$, we have $V_{1}([a,1])\geq\frac{2}{n}V_{1}([0,1])=\frac{2}{n}$. This implies $a\leq\frac{n-2}{n}$ and $b=\frac{a+1}{2}\leq\frac{n-1}{n}$. On the other hand, by reporting $v_{1}^{\prime}$, agent $a_{1}$ can then get the piece $a,b^{\prime}]$ with $b^{\prime}\in[1-2\varepsilon,1-\varepsilon]$. We see that $b^{\prime}>b$. Thus, the manipulation is safe and profitable. ∎
Theorem 6.9.
The RAT-degree of Bu, Song, and Tao’s moving knife procedure is $1$.
Proof.
It was proved in BU2023Rat that the mechanism is not 0-known-agent safely-manipulable. The proof that it is $1$-known-agents safely-manipulable is similar to the proof for Ortega and Segal-Halevi’s moving knife procedure, with the same $v_{1},v_{1}^{\prime}$ and $v_{2}$. It suffices to notice that the first n2n-2 equal-division-points are the same for $v_{1}$ and $v_{1}^{\prime}$, where the last equal-division-point of $v_{1}^{\prime}$ is to the right of $v_{1}$’s. Given that agent $a_{2}$ will always receive a piece after agent $a_{1}$, the same analysis in the previous proof can show that the manipulation is safe and profitable. ∎
6.4. Volatile Priority Cake Cutting: Proportional, Pareto-Optimal, with RAT-degree n1n-1
In this section, we provide a mechanism with RAT-degree n1n-1 that always outputs proportional and Pareto-optimal allocations. The mechanism uses some similar ideas as the one in Section 5.5.
Here, the given priority function $\Gamma$, still takes as input a valuation profile $(v_{1},\ldots,v_{n})$ but outputs an order $\pi$ of the $n$ agents. In the appendix, we formally adapt the volatility property (defined in Section 5.5) to the cake-cutting setting and prove that such function can be constructed.
Proposition 6.10.
There exists a volatile priority rule $\Gamma$ for cake cutting.
Let $\mathcal{X}^{\mathrm{PROP}}$ be the set of all proportional allocations and let $\pi_{i}$ be the $i$-th agent in the order $\pi$. Then, the mechanism outputs an allocation in $\mathcal{X}^{\mathrm{PROP}}$ in the following “leximax” way:
  1. (1)
    the allocation maximizes agent $\pi_{1}$’s utility;
  2. (2)
    subject to (1), the allocation maximizes agent $\pi_{2}$’s utility;
  3. (3)
    subject to (1) and (2), the allocation maximizes agent $\pi_{3}$’s utility;
  4. (4)
    $\cdots$
By definition, the mechanism always outputs a proportional allocation. It is straightforward to check that it outputs a Pareto-efficient allocation. We further prove that:
Theorem 6.11.
thm:volatile-cake-degree-n-1 The RAT-degree of the Volatile Priority Cake Cutting is n1n-1.
The proof is similar to the one for indivisible goods (LABEL:thm:volatile-good-alloc-degree-n-1) and is provided in the appendix. Finally, we analyze the run-time of our mechanism.
Theorem 6.12.
The Volatile Priority Cake Cutting mechanism can be computed in polynomial time.
Proof sketch.
Once $\pi$ is determined, the allocation is computed by solving a sequence of linear programs. Let $x_{it}$ be the length of the $t$-th uniform segment allocated to agent $a_{i}$. Then an agent $a_{i}$’s utility is a linear expression $\sum_{t=1}^{m}v_{i}(X_{t})x_{it}$, and requiring an agent’s utility is at least some value (e.g., her proportional share) is a linear constraint. ∎
7. Single-Winner Ranked Voting
We consider $n$ voters (the agents) who need to elect one winner from a set $C$ of $m$ candidates. The agents’ preferences are given by strict linear orderings $\succ_{i}$ over the candidates.
When there are only two candidates, the majority rules and its variants (weighted majority rules) are truthful. With three or more candidates, the Gibbard–Satterthwaite Theorem (gibbard1973manipulation; satterthwaite1975strategy) implies that the only truthful rules are dictatorships. Our goal is to find non-dictatorial rules with a high RAT-degree.
We focus on positional voting rules, which parameterized by a vector of scores, $\mathbf{s}=(s_{1},\ldots,s_{m})$, where $s_{1}\leq\cdots\leq s_{m}$ and $s_{1}<s_{m}$. Each voter reports his entire ranking of the $m$ candidates. Each such ranking is translated to an assignment of a score to each candidate: the lowest-ranked candidate is given a score of $s_{1}$, the second-lowest candidate is given $s_{2}$, etc., and the highest-ranked candidate is given a score of $s_{m}$. The total score of each candidate is the sum of scores he received from the rankings of all $n$ voters. The winner is the candidate with the highest total score.
Formally, for any subset $N^{\prime}\subseteq N$ and any candidate $c\in C$, we denote by $\operatorname{score}_{N^{\prime}}(c)$ the total score that $c$ receives from the votes of the agents in $N^{\prime}$. Then the winner is $\arg\max_{c\in C}\operatorname{score}_{N}(c)$. If there are several agents with the same maximum score, then the outcome is considered a tie.
Common special cases of positional voting are plurality voting, in which $\mathbf{s}=(0,0,0,\ldots,0,1)$, and anti-plurality voting, in which $\mathbf{s}=(0,1,1,\ldots,1,1)$. By the Gibbard–Satterthwaite theorem, all positional voting rules are manipulable, so their RAT-degree is smaller than $n$. But, as we will show next, some positional rules have a higher RAT-degree than others.
Results.
conitzer2011dominating showed that all positional voting rules are immune to dominating manipulations when $n\geq 6(m-2)$ — in our terms, that they have positive RAT-degree. We extend their results for the case where $n\geq 2m$ by proving that all positional voting rules have an RAT-degree between $\approx n/m$ and $\approx n/2$. These bounds are almost tight: the upper bound is attained by plurality and the lower bound is attained by anti-plurality (up to small additive constants).
These results raise the question of whether some other, non-positional voting rules have RAT-degrees substantially higher than $n/2$. Using our volatile priority approach (see Section 5), we could choose a “dictator”, and take her first choice. This deterministic mechanism has RAT-degree n1n-1, any manipulation risks losing the chance to be the dictator. However, besides the fact that this is an unnatural mechanism, it suffers from other problems such as the no-show paradox (a participating voter might affect the selection rule in a way that will make another agent a dictator, which might be worse than not participating at all).
Our main open problem is therefore to devise natural voting rules with a high RAT-degree.
Open Question 7.1.
Does there exist a non-dictatorial voting rule that satisfies the participation criterion (i.e. does not suffer from the no-show paradox), with RAT-degree larger than $\left\lceil n/2\right\rceil+1$?
conitzer2011dominating further proved, in our terms, that any Condorcet-consistent voting rule has a positive RAT-degree. This raises the following question
Open Question 7.2.
What are the exact RAT-degrees of known Condorcet-consistent rules?
Moulin1982vetoPower suggested the notion of veto power for voting. We provide further details on the relationship between the two notions in Section 7.3.
Mechanism RAT-Degree
Lower Bound Upper Bound
Positional Voting Rules (assuming $n\geq 2m$) $\left\lfloor(n+1)/m\right\rfloor-1$ $\left\lceil n/2\right\rceil+1$
Plurality (assuming $n\geq 5$) $\left\lfloor n/2\right\rfloor+1$ $\left\lceil n/2\right\rceil+1$
Anti-Plurality (assuming $n\geq m^{2}$) $\left\lfloor(n+1)/m\right\rfloor-1$ $\left\lfloor n/m\right\rfloor+1$
Table 5. Single-Winner Ranked Voting with $m\geq 3$: Summary of Results.
Throughout the analysis, we consider a specific agent Alice, who looks for a safe profitable manipulation. Her true ranking is $c_{m}\succ_{A}\cdots\succ_{A}c_{1}$. We assume that, for any $j>i$, Alice strictly prefers a victory of $c_{j}$ to a tie between $c_{j}$ and $c_{i}$, and strictly prefers this tie to a victory of $c_{i}$. [^4] We could also assume that ties are broken at random, but this would require us to define preferences on lotteries, which we prefer to avoid in this paper.
7.1. Positional Voting Rules: General Bounds
In the upcoming lemmas, we identify the manipulations that are safe and profitable for Alice under various conditions on the score vector $\mathbf{s}$. We assume throughout that there are $m\geq 3$ candidates, and that $n$ is sufficiently large. We allow an agent to abstain, which means that his vote gives the same score to all candidates. [^5] We need the option to abstain in order to avoid having different constructions for even $n$ and odd $n$; see the proofs for details.
The first lemma implies a lower bound of $\approx n/m$ on the RAT-degree of positional voting rules.
Lemma 7.3.
lem:lower-positional In any positional voting rule for $m\geq 3$ candidates, if the number of known agents is at most (n+1)/m2(n+1)/m-2, then Alice has no safe profitable manipulation.
Proof sketch.
For any combination of rankings of the known agents, it is possible that the unknown agents vote in a way that balances out the votes of the known agents, such that all agents have almost the same score; if Alice is truthful, there is a tie between two candidates, and if she manipulates, the worse of these candidates win. ∎
We now prove an upper bound of $\approx n/2$ on the RAT-degree. We need several lemmas.
Lemma 7.4.
lem:s2¿s1 Let $m\geq 3$ and $n\geq 2m$. If $s_{2}>s_{1}$ and there are $k\geq\left\lceil n/2\right\rceil+1$ known agents, then switching the bottom two candidates ($c_{2}$ and $c_{1}$) may be a safe profitable manipulation for Alice.
Proof sketch.
For some votes by the known agents, $c_{1}$ has no chance to win, so the worst candidate for Alice that could win is $c_{2}$. Therefore, switching $c_{1}$ and $c_{2}$ cannot harm, but may help a better candidate win over $c_{2}$. ∎
LABEL:lem:s2>s1 can be generalized as follows.
Lemma 7.5.
lem:st1¿st Let $m\geq 3$ and $n\geq 2m$. For every integer $t\in\{1,\ldots,m-2\}$, if $s_{t+1}>s_{t}=\cdots=s_{1}$, and there are $k\geq\left\lceil n/2\right\rceil+1$ known agents, then switching $c_{t+1}$ and $c_{t}$ may be a safe profitable manipulation.
Proof sketch.
For some votes by the known agents, all candidates $c_{1},\ldots,c_{t}$ have no chance to win, so the worst candidate for Alice that could win is $c_{t+1}$. Therefore, switching $c_{t}$ and $c_{t+1}$ cannot harm, but can help better candidates win over $c_{t+1}$. ∎
Lemma 7.6.
lem:sm¿sm1 Let $m\geq 3$ and $n\geq 4$. If $s_{m}>s_{m-1}$ and there are $k\geq\left\lceil n/2\right\rceil+1$ known agents, then switching the top two candidates ($c_{m}$ and $c_{m-1}$) may be a safe profitable manipulation for Alice.
Proof sketch.
For some votes by the known agents, the manipulation is safe since $c_{m}$ has no chance to win, and it is profitable as it may help $c_{m-1}$ win over worse candidates. ∎
Combining the lemmas leads to the following bounds on the RAT-degree:
Theorem 7.7.
For any positional voting rule with $m\geq 3$ candidates:
(a) The RAT-degree is at least $\left\lfloor(n+1)/m\right\rfloor-1$;
(b) When $n\geq 2m$, the RAT-degree is at most $\left\lceil n/2\right\rceil+1$.
Proof.
The lower bound follows immediately from LABEL:lem:lower-positional.
For the upper bound, consider a positional voting rule with score-vector $\mathbf{s}$. Let $t\in\{1,\ldots,m-1\}$ be the smallest index for which $s_{t+1}>s_{t}$ (there must be such an index by definition of a score-vector).
If $t\leq m+2$, then LABEL:lem:st1>st implies that, for some votes by the $\left\lceil n/2\right\rceil+1$ known agents, switching $c_{t+1}$ and $c_{t}$ may be a safe and profitable manipulation for Alice.
Otherwise, t=m1t=m-1, and LABEL:lem:sm>sm1 implies the same.
In all cases, Alice has a safe profitable manipulation. ∎
7.2. Plurality and Anti-plurality
We now show that the bounds of Theorem 7.7 are tight up to small additive constants.
We first show that the upper bound of $\approx n/2$ is attained by the plurality voting rule, which is the positional voting rule with score-vector $(0,0,0,\ldots,0,1)$.
Lemma 7.8.
lem:lower-plurality In the plurality voting rule with $n\geq 5$ agents, if the number of known agents is at most $n/2$, then Alice has no safe profitable manipulation.
Proof sketch.
When there are at most $n/2$ known agents, there are at least n/21n/2-1 unknown agents. For some votes of these unknown agents, the outcome when Alice votes truthfully is a tie between Alice’s top candidate and another candidate. But when Alice manipulates, the other candidate wins. ∎
Combining LABEL:lem:sm>sm1 and LABEL:lem:lower-plurality gives an almost exact RAT-degree of plurality voting.
Theorem 7.9.
With $m\geq 3$ candidates and $n\geq 5$ agents, the RAT-degree of plurality voting is $n/2+1$ when $n$ is even; it is between $\left\lfloor n/2\right\rfloor+1$ and $\left\lceil n/2\right\rceil+1$ when $n$ is odd.
Next, we show that the lower bound of $\approx n/m$ is attained by the anti-plurality voting rule.
Lemma 7.10.
With $m\geq 3$ and $n\geq m^{2}$, the RAT-degree of anti-plurality is at most $\lfloor n/m\rfloor+1$.
Proof sketch.
We show that switching the two least-favorite candidates ($c_{1}$ and $c_{2}$) constitutes a safe and profitable manipulation for Alice, as the number of known agents is sufficient to ensure that her least-preferred candidate cannot win. ∎
To this end, we also prove several upper bounds on the RAT-degree for more general score-vectors.
The following lemma strengthens LABEL:lem:s2>s1.
Lemma 7.11.
lem:z:s2¿s1 Let $\ell\in\{2,\ldots,m-1\}$ be an integer. Consider a positional voting setting with $m\geq 3$ candidates and $n\geq(\ell+1)m$ agents. Denote $s_{\mathrm{top:}\ell}:=\sum_{j=m-\ell+1}^{m}s_{j}=$ the sum of the $\ell$ highest scores and $s_{\mathrm{bot:}\ell}:=\sum_{j=1}^{\ell}s_{j}=$ the sum of the $\ell$ lowest scores.
If $s_{2}>s_{1}$ and there are $k$ known agents, where
$k>\frac{\ell s_{m}-s_{\mathrm{bot:}\ell}}{\ell s_{m}+s_{\mathrm{top:}\ell}-s_{\mathrm{bot:}\ell}-\ell s_{1}}n,$
then switching the bottom two candidates ($c_{2}$ and $c_{1}$) may be a safe profitable manipulation for Alice.
Proof sketch.
The proof has a similar structure to that of LABEL:lem:s2>s1. Note that the expression at the right-hand side can be as small as $\frac{1}{\ell+1}n$ (for the anti-plurality rule), which is much smaller than the $\left\lceil n/2\right\rceil+1$ known agents required in LABEL:lem:s2>s1. Still, we can prove that, for some reports of the known agents, the score of $c_{1}$ is necessarily lower than the arithmetic mean of the scores of the $\ell$ candidates $\{c_{m},c_{2},\cdots,c_{\ell}\}$. Hence, it is lower than at least one of these scores. Therefore $c_{1}$ still cannot win, so switching $c_{1}$ and $c_{2}$ is safe. ∎
In particular, for the anti-plurality rule the condition in LABEL:lem:z:s2>s1 for $\ell=m-1$ is $k>n/m$, which implies a lower bound of $\left\lfloor n/m\right\rfloor+1$. Combined with the general upper bound of Theorem 7.7, we get:
Theorem 7.12.
With $m\geq 3$ candidates and $n\geq m^{2}$ agents, The RAT-degree of anti-plurality voting is at least $\left\lfloor(n+1)/m\right\rfloor-1$ and at most $\left\lfloor n/m\right\rfloor+1$.
Intuitively, the reason that anti-plurality fares worse than plurality is that, even with a small number of known agents, it is possible to deduce that some candidate has no chance to win, and therefore there is a safe manipulation.
While we do not yet have a complete characterization of the RAT-degree of positional voting rules, our current results already show the strategic importance of the choice of scores.
7.3. RAT-Degree and Veto Power
In this section, we examine the relation between the RAT-degree and the notion of veto power introduced by Moulin1982vetoPower. Formally, a coalition of voters of size $t$ is said to have veto power $v(t)$ if it can prevent any set of $v(t)$ candidates from winning, regardless of the preferences of the remaining voters.
Initially, we thought that the veto power can provide an upper bound on the RAT-degree, as a manipulator who knows that some candidate has no chance of winning, can use this knowledge to obtain a safe and profitable manipulation. Therefore, we expected that the minimal size of a coalition with positive veto power would be an upper bound on the RAT-degree. However, this claim appears to fail in general:
Lemma 7.13.
There exist voting rules for which, even when the manipulator knows that one candidate has no chance of winning, this is not sufficient to ensure the existence of a safe and profitable manipulation.
Proof.
Without loss of generality, assume $m$ is even and consider the m/2m/2–approval rule. Under m/2m/2–approval, each voter gives score $1$ to his top $m/2$ candidates and 0 to the rest; thus any unilateral change by $i$ can change every candidate’s total score by at most $\pm 1$.
Let Alice be the manipulator with true preferences $c_{m}\succ m_{m-1}\succ\cdots\succ c_{1}$. We show that if Alice only knows that one candidate has no chance of winning, any profitable manipulation is not safe. Consider the following cases:
  • The eliminated candidate is one of Alice’s top $m/2$ candidates:
    In this case, the manipulations that are potentially profitable are to move this eliminated candidate from the approved set (score $1$) to the unapproved set (score 0), and to promote instead one of the unapproved candidates to the approved set. Any other manipulation contributes exactly the same total scores as truthful reporting and therefore cannot be profitable. However, this swap may cause a worse outcome for Alice: it is possible that when she tell the truth the result is a tie between the her favorite candidate and the less-preferred candidate that she promoted and so when she manipulates the less-preferred candidate wins. Therefore, such a manipulation is not safe.
  • The eliminated candidate is among Alice’s bottom $m/2$ candidates:
    Here, the only potentially profitable manipulation is to promote this eliminated candidate into the approved set (score $1$) and to demote the least-preferred among the currently approved candidates. Any other manipulation yields the same contribution as truth-telling. While this move might seem harmless — since the promoted candidate cannot win — it may reduce the score of one of the approved candidates enough to let a less-preferred candidate win, which is therefore not safe.
This lemma shows that knowing that a single candidate has no chance of winning is generally insufficient to guarantee a safe and profitable manipulation. However, we show that such knowledge can still be helpful in some cases. In particular, under certain voting rules, knowing that several candidates are ruled out might help, where the required number of candidates depends on the structure of the specific rule. We illustrate this through two partial results that hold for any positional voting rule. We prove that knowing that some candidates are ruled out may enable a manipulation that is safe; however, it is unclear whether it is necessarily profitable. We leave this as an open question and provide full details below. Specifically, Lemma 7.14 establishes a connection based on the first position where the score strictly increases, while Lemma 7.15 relates the RAT-degree to the length of the largest sequence of identical scores.
Lemma 7.14.
Consider any position voting rule with $s_{1}<s_{m}$. Let $\ell\in\{1,\ldots,m-1\}$ be the smallest position such that $s_{\ell}<s_{\ell+1}$.  Let $t$ be the smallest coalition size with veto power at least $\ell$— i.e., $v(t)\geq\ell$. Then, if the number of known-agents is at least $t$, then the manipulator has safe (but not necessarily profitable) manipulation.
Proof.
Let Alice be the manipulator with true preferences $c_{m}\succ m_{m-1}\succ\cdots\succ c_{1}$. Let $K$ be a set of $t$ voters with veto power $\ell$. By definition, they can rule out any set of $\ell$ candidates. Consider the case where they rule out Alice’s $\ell$ least-preferred candidates.
Notice that $c_{\ell}$ and $c_{\ell+1}$ are the candidate at position $\ell$ and $\ell+1$ in Alice’s truthful preferences, respectively. Consider the following manipulation for Alice: switching between $c_{\ell}$ and $c_{\ell+1}$.
First, as $s_{\ell}<s_{\ell+1}$, swapping $c_{\ell}$ and $c_{\ell+1}$ strictly increases $c_{\ell}$’s score and strictly decreases $c_{\ell+1}$’s score. This means that the manipulation is safe: promoting $c_{\ell}$ cannot make this candidate win, since it is one of the candidates ruled out by $K$.
The manipulation might be profitable as the competition is only between $c_{\ell+1}$ and candidates that are more preferred by Alice, so decreasing the score of $c_{\ell+1}$ can only benefit her. However, for a manipulation to be profitable, we must show that there exists some profile, consistent with the agent’s knowledge, under which the outcome is strictly improved compared to truthful reporting. In particular, if we could guarantee, for example, the existence of a profile in which there is a tie between $c_{\ell+1}$ and a more preferred candidate, then the manipulation would break the tie in favor of the latter, yielding a strictly better outcome. However, since we only know that certain candidates are ruled out by the agent’s knowledge, it is unclear whether such a profile must necessarily exist [^6] Assume, for example, that the only profile for the known-agents that rules out the desired candidates (Alice’s $\ell$ least-favorite candidates) also rules out candidate $c_{\ell+1}$. In this case, $c_{\ell+1}$ has no chance of winning and therefore the manipulation is not profitable.. ∎
Lemma 7.15.
Consider any position voting rule with $s_{1}<s_{m}$. Let $q\{1,\ldots,m-1\}$ be the length of the largest sequence of identical scores:
$q=\max_{1\leq j\leq m-1}\{\,x\mid s_{j}=s_{j+1}=\cdots=s_{j+x-1}\,\}.$
Let $t$ be the smallest coalition size with veto power at least mqm-q — i.e., $v(t)\geq m-q$.
Then, if the number of known-agents is at least $t$, then the manipulator has safe (but not necessarily profitable) manipulation.
Proof.
Let Alice be the manipulator with true preferences $c_{m}\succ c_{m-1}\succ\cdots\succ c_{1}$. Let $K$ be a set of $t$ voters with veto power mqm-q. By definition, they can rule out any set of mqm-q candidates. Consider the case where they rule out all but a set of $q$ candidates corresponding to the largest equal-score sequence (if there are several such sets, pick one). Formally, let $j$ be the smallest index of that sequence and $C_{q}:=\{c_{j},\ldots,c_{j+q-1}\}$; then the known agents rule out all candidates that are not in $C_{q}$.
Since $s_{1}<s_{m}$, at least one of the two cases must hold:
  • $j>1$:
    Consider the following manipulation for Alice: switching between $c_{1}$ and $c_{j}$. Notice that $s_{j}>s_{1}$ (by the minimality of the length of the equal-score sequence), thus the manipulation strictly increases $c_{1}$’s score and strictly decreases $c_{j}$’s score.
    As the competition is only between the candidates in $C_{q}$, increasing $c_{1}$’s score cannot harm Alice (as $c_{1}$ has no chance of winning), while decreasing $c_{j}$, which is the least-preferred candidate in $C_{q}$, can only benefit her. Therefore, the manipulation is safe.
    It might also be profitable: if under Alice’s truthful report the outcome is a tie between $c_{j}$ and a more-preferred candidate, then this manipulation breaks the tie in favor of her more preferred one. It is unclear if such a profile must necessarily exist.
  • j+q1<mj+q-1<m:
    In this case, switching between $c_{j+q-1}$ and $c_{m}$ is safe and profitable. Here, we can conclude that $s_{j+q-1}<s_{m}$. which means that the manipulation strictly increases $c_{j+q-1}$’s score and strictly decreases $c_{m}$’s score.
    As the competition is only between the candidates in $C_{q}$, decreasing $c_{m}$’s score cannot harm Alice (as it has no chance of winning). However, increasing $c_{j+q-1}$, which is her most-preferred candidate in $C_{q}$, can only benefit her. Thus, the manipulation is safe.
    The manipulation might be profitable: if under her truthful report the outcome is a tie between $c_{j+q-1}$ and a least-preferred candidate, then this manipulation breaks the tie in favor of her more preferred one. Unfortunately, it is unclear whether such a profile must necessarily exist.
This raises the following question:
Open Question 7.16.
Is there a direct relation between Moulin’s veto power and the RAT degree, and if so, how can it be characterized?
8. Two-sided Matching
In this section, we consider mechanisms for two-sided matching. Here, the $n$ agents are divided into two disjoint subsets, $M$ and $W$, that need to be matched to each other. The most common examples are men and women or students and universities. Each agent has a strict preference order over the agents in the other set and being unmatched – for each $m\in M$, an order $\succ_{m}$ over $W\cup\{\phi\}$; and for each $w\in W$ an order, $\succ_{w}$, over $M\cup\{\phi\}$.
A matching between $M$ to $W$ is a mapping $\mu$ from $M\cup W$ to $M\cup W\cup\{\phi\}$ such that (1) $\mu(m)\in W\cup\{\phi\}$ for each $m\in M$, (2) $\mu(w)\in M\cup\{\phi\}$ for each $w\in W$, and (3) $\mu(m)=w$ if and only if $\mu(w)=m$ for any $(m,w)\in M\times W$. When $\mu(a)=\phi$ it means that agent $a$ is unmatched under $\mu$.
A mechanism in this context gets the preference orders of all agents and returns a matching. See gonczarowski2024structural for a recent description of the structure of matching mechanisms.
Results.
Our results for this problem are preliminary, so we provide only a brief overview here, with full descriptions and proofs in the appendix. We believe, however, that this is an important problem and that our new definition opens the door to many interesting questions.
We start with the well-known deferred acceptance mechanism of gale1962college, which always returns a stable matching—no agent prefers being unmatched over their assigned match, and there is no pair of agents who would both prefer to be matched to each other over their current assignments. It is known that no stable matching mechanism is truthful for all agents (roth1982economics). Indeed, deferred acceptance is known to be truthful only for one side of the market. But what happens on the other side? Our analysis reveals that the RAT-degree of deferred acceptance is very low: it is at least $1$ and at most $3$. The proof of the upper bound relies on truncation, where an agent falsely reports preferring to remain unmatched over certain options. We further show that even when agents are required to report complete preferences—thus ruling out this type of manipulation—the RAT-degree is at most $5$.
This raises the following important and interesting question:
Open Question 8.1.
Is there a stable matching mechanism with RAT-degree in $\Omega(n)$?
We also examine the Boston mechanism (abdulkadirouglu2003school), which is a widely used in practice for assigning students to schools. We establish an upper bound of $2$ on its RAT-degree.
Mechanism RAT-Degree Properties
Lower Bound Upper Bound
Deferred Acceptance (DA) 11 33 Stable, Truthful for $M$
DA under Complete Preferences 11 55
Boston 0 2
Table 6. Matchings: Summary of Results.
8.1. Deferred Acceptance (Gale-Shapley)
The deferred acceptance algorithm (gale1962college) is one of the most well-known mechanisms for computing a stable matching. In this algorithm, one side of the market — here, $M$ — proposes, while the other side — $W$ — accepts or rejects offers iteratively. For completeness, we provide its description in the appendix.
It is well known that the mechanism is truthful for the proposing side ($M$) but untruthful for the other side ($W$). That is, the agents in $W$ may have an incentive to misreport their preferences to obtain a better match. Moreover, there is provably no mechanism for two-sided matching that is truthful for both sides (roth1982economics).
This section provides a more nuanced analysis of the amount of knowledge required by $W$ agents to manipulate safely. We focus on a specific agent $w_{1}\in W$, with ranking $m_{1}\succ_{w_{1}}m_{2}\succ_{w_{1}}\cdots$. There are three kinds of potential manipulations for any $w_{1}\in W$:
  1. (1)
    Demoting some $m_{j}\in M$ from above $\phi$ to below $\phi$ (i.e., claiming that an acceptable partner is unacceptable for her). This is equivalent to reporting only a prefix of the ranking, sometimes called truncation (roth1999truncation; ehlers2008truncation; coles2014optimal).
  2. (2)
    Promoting some $m_{j}\in M$ from below $\phi$ to above $\phi$ (i.e., claiming that an unacceptable partner is acceptable for her).
  3. (3)
    Reordering some $m_{i},m_{j}\in M$ (i.e., claiming that she prefers $m_{i}$ to $m_{j}$ where in fact she prefers $m_{j}$ to $m_{i}$).
Lemma 8.2.
lem:da-truncation In Deferred Acceptance with $k\geq 3$ known agents, there may be a safe and profitable truncation manipulation for $w_{1}$.
In some settings, it is reasonable to assume that agents always prefer being matched if possible. In such cases, the mechanism is designed to accept only preferences over agents from the opposite set (or equivalently, orders where being unmatched is always the least preferred option). Clearly, under this restriction, truncation is not a possible manipulation. We prove that even when truncation is not possible, the RAT-degree is bounded.
Lemma 8.3.
lem:da-switch In Deferred Acceptance without truncation, with $k\geq 5$ known agents, there may be a safe and profitable reorder manipulation for $w_{1}$.
Remark 8.4.
In the women-proposing variant, $w_{1}$ gets $m_{1}$ both when she is truthful and when she manipulates.
Hence, A ”mixture” variant of DA, which mixes between men-proposing and women-proposing, will not have a higher RAT degree.
We conjecture that the numbers $3$ and $5$ are tight, but currently have only a weaker lower bound of $1$, which follows from the following lemma:
Lemma 8.5.
In Deferred Acceptance, no agent has a safe manipulation.
Combining the above lemmas gives:
Theorem 8.6.
The RAT-degree of Deferred Acceptance is at least $1$.
It is at most $3$ when truncation is allowed, and at most $5$ when truncation is not allowed.
8.2. Boston Mechanism
The Boston mechanism (abdulkadirouglu2003school) is a widely used mechanism for assigning students or schools. It is not truthful for both sides. Moreover, we show that it can be safely manipulated with little information.
Lemma 8.7.
lem:prop-boston In the Boston Mechanism with $k\geq 2$ known agents, there may be a safe and profitable reorder manipulation for some $m_{1}\in M$.
The lemma implies:
Theorem 8.8.
The RAT-degree of the Boston mechanism is at most $2$.
9. Discussion and Future Work
Our main goal in this paper is to encourage a more quantitative approach to truthfulness that can be applied to various problems. When truthfulness is incompatible with other desirable properties, we aim to find mechanisms that are “as hard to manipulate as possible”, where hardness is measured by the amount of knowledge required for a safe manipulation.
Avoiding Risk.
Our definition applies only to agents who entirely avoid even a small risk. This definition can be justified based on the well-known behavioral phenomenon called ”zero-risk bias”(crosby2021laws; wells2017monetizing). For example, experiments reported in the paper ’Prospect Theory: An Analysis of Decision under Risk’, by kahneman2013prospect, indicates that people underweight outcomes that are merely probable, in comparison with outcomes that are obtained with certainty.
Another justification stems from our distinction between two types of risk: the risk due to the randomization of the mechanism, and the risk due to uncertainty about other agents’ preferences. Our risk-avoidance assumption concerns only the second type of risk, for which there is usually no known probability distribution. Without such knowledge, agents are facing a non-quantifiable unknown risk (often referred to as ambiguity). [^7] This is related to another behavioral phenomenon known as ”ambiguity aversion” or ”uncertainty aversion”. We believe that it is more reasonable to assume agents avoid this unknown risk altogether as they cannot even reliably distinguish between high and low risks. To compute expected utility with respect to this risk, one would need to adopt a Bayesian framework, which often requires strong and somewhat artificial assumptions about the strategies of other agents. We leave this direction for future work.
Randomized Mechanisms.
This distinction between the two types of risk naturally leads to the topic of randomized mechanisms. In some settings, RAT-degree is more interesting for deterministic mechanisms, as many impossibility results apply only to deterministic mechanisms. However, in contexts such as voting, where there are impossibility results for randomized mechanisms too, the RAT-degree is applicable and potentially useful. In such settings, one can assume that agents avoid the non-quantifiable unknown risk that comes from the uncertainty about other agents’ preferences, while treating the known risk induced by the mechanism construction using standard risk models—such as risk-neutral or risk-averse—aimed at maximizing expected utility.
Utility: Zero vs. Positive.
Our model assumes that even a small positive gain can have a significant impact on behavior. For instance, the RAT-degree of the first-price auction is 0, but increases to $1$ when we introduce a fixed positive discount, even when the discount is arbitrarily small (see Section 4). This assumption is supported by findings in behavioral economics. For example, the paper ’Zero as a Special Price: The True Value of Free Products’, by shampanier2007zero, shows that behavior changes significantly when something is free versus when it carries even a tiny cost. A related real-world example is the shift in behavior following the introduction of a small charge on plastic bags in many countries. When bags were free, most people used them without thinking; even a tiny positive cost was sufficient for ‘nudging’ people to bring reusable bags.
Tradeoffs.
A particularly interesting future direction is to explore the tradeoffs between a high RAT-degree and other desirable properties, such as fairness. For example, in the case of two-sided matching (Section 8), we know that stability can be achieved with a low RAT-degree of at most $3$, but is impossible to achieve with a RAT-degree of $n$. Where exactly does the boundary lie? Can we characterize the RAT-degree in relation to different fairness properties?
Beyond Worst-Case.
We defined the RAT-degree as a “worst case” concept: to prove an upper bound, we find a single example of a safe manipulation. This is similar to the situation with the classic truthfulness notion, where to prove non-truthfulness, it is sufficient to find a single example of a manipulation. To go beyond the worst case, one could follow relaxations of truthfulness, such as truthful-in-expectation (lavi2011truthful) or strategyproofness-in-the-large (azevedo2019strategy), and define similarly “RAT-degree in expectation” or “RAT-degree in the large”.
Information About the Known-Agents.
Our definition assumes that whenever an agent is known, we know their exact preference—that is, the precise preferences $P_{i}$ they will report from their domain $\mathcal{D}_{i}$. In practice, however, we often have only partial information on the known-agents; for example, we might know that their preferences belong to a smaller subset of their domain. Consider Table 1 in Section 3.1. Such information still allows us to consider only a strict subset of the columns, leading to a weaker requirement than standard truthfulness. However, our results show that even under the assumption that we know the exact preferences, identifying the RAT-degree can already be highly nontrivial. For this reason, we only highlight this direction as a promising avenue for future research.
Changing Quantifiers.
One could argue for a stronger definition requiring that a safe manipulation exists for every possible set of $k$ known-agents, rather than for some set, or similarly for every possible preference profile for the known agents rather than just in some profile. However, we believe such definitions would be less informative, as in many cases, a manipulation that is possible for some set of $k$ known-agents, is not possible for any such set. For example, in the first-price auction with discount (see Section 4.2), the RAT-degree is $1$ under our definition. But if we required the the knowledge on any agent’s bid would allow manipulation rather than just one, the degree would automatically jump to n1n-1, making the measure far less meaningful.
Combining the ”known agents” concept with other notions.
We believe that the “known agents” approach can be used to quantify the degree to which a mechanism is robust to other types of manipulations (besides safe manipulations), such as “always-profitable” manipulations or “obvious” manipulation. Accordingly, one can define the “max-min-strategyproofness degree” or the “NOM degree” (see the ).
Alternative Information Measurements.
Another avenue for future work is to study other ways to quantify truthfulness. For example, instead of counting the number of known agents, one could count the number of bits that an agent should know about other agents’ preferences in order to have a safe manipulation. The disadvantage of this approach is that different domains have different input formats, and therefore it would be hard to compare numbers of bits in different domains (see the   for more details).
Other applications.
The RAT-degree can potentially be useful in any social-choice setting in which truthful mechanisms are not known, or lack other desirable properties. Examples include combinatorial auctions, multiwinner voting, budget aggregation and facility location.
10. Acknowledgement
This research is partly supported by the Israel Science Foundation grants 712/20, 3007/24, 2697/22 and 1092/24. The research of Biaoshuai Tao is supported by the National Natural Science Foundation of China (No. 62472271) and the Key Laboratory of Interdisciplinary Research of Computation and Economics (Shanghai University of Finance and Economics), Ministry of Education.
We sincerely appreciate Alexandros Psomas for his valuable contribution to this work. We are grateful to Sophie Bade, Avinatan Hassidim, Assaf Romm, Yonatan Aumann, Vincent Conitzer, Jerome Lang, and Reshef Meir for their insights and helpful answers, to Paritosh Verma for his contribution; and to Michael Greinecker for his helpful responses. [^8] https://economics.stackexchange.com/q/59899/385 and https://economics.stackexchange.com/a/24369/385
Lastly, we would like to thank the reviewers in EC 2025 and COMSOC 2025 for their helpful comments.
Appendix A Open Questions: Summary
  • Indivisible Goods Allocations
    • 5.1: Is there a polynomial time EF1 goods allocation rule with RAT-degree n1n-1?
    • 5.2: What is the RAT-degree of The Envy-Cycle Elimination Mechanism by lipton2004approximately?
    • 5.17: What is the RAT-degree of Volatile Priority Round Robin?
    • 5.12: Are there any variants of MNW with a RAT degree larger than 1?
  • Cake Cutting
    • 6.1:Is there a proportional connected cake-cutting rule with RAT-degree at least $2$?
    • 6.2: What is the RAT-degree of the Even-Paz algorithm?
    • 6.3: What is the RAT-degree of the maximum Nash welfare mechanism?
  • Single-Winner Voting
    • 7.1: Does there exist a non-dictatorial voting rule that satisfies the participation criterion (i.e. does not suffer from the no-show paradox), with RAT-degree larger than $\left\lceil n/2\right\rceil+1$?
    • 7.2: What are the exact RAT-degrees of known Condorcet-consistent rules?
    • 7.16: Is there a direct relation between Moulin’s veto power and the RAT degree, and if so, how can it be characterized?
  • Two-Sided Matching
    • 8.1: Is there a stable matching mechanism with RAT-degree in $\Omega(n)$?
Appendix B Related Work: Extended Discussion
To maintain consistency with the proceedings version, we have kept the appendix, even though its content is now integrated into the main body of the paper. For the full discussion, please refer to the Related Work section in this version.